### v1

# Stochastic dominance for super heavy-tailed random variables

Abstract
We introduce a class of super heavy-tailed distributions and establish the inequality that any weighted average of independent and identically distributed super heavy-tailed random variables stochastically dominates one such random variable. We show that many commonly used extremely heavy-tailed (i.e., infinite-mean) distributions, such as the Pareto, Fréchet, and Burr distributions, belong to the class of super heavy-tailed distributions. The established stochastic dominance relation is further generalized to allow negatively dependent or non-identically distributed random variables. In particular, the weighted average of non-identically distributed random variables stochastically dominates their distribution mixtures. Applications of these results in portfolio diversification, goods bundling, and inventory management are discussed. Remarkably, in the presence of super heavy-tailedness, the results that hold for finite-mean models in these applications are flipped.
Keywords: heavy-tailed distributions; stochastic order; negative dependence; infinite mean.
1 Introduction
Distributions with infinite mean are ubiquitous in the realm of banking and insurance, and they are particularly useful in modeling catastrophic losses (Ibragimov et al. (2009)), operational losses (Moscadelli, (2004)), costs of cyber risk events (Eling and Wirfs, (2019)), and financial returns from technology innovations (Silverberg and Verspagen, (2007)); see also Chen et al. (2024b) for a list of empirical examples of distributions with infinite mean.
As the world is arguably finite (e.g., any loss is bounded by the total wealth in the world), why should we use models with infinite mean as mathematical tools? There are perhaps two reasons. The first reason is that, after careful statistical analysis, infinite-mean models often fit extremely heavy-tailed datasets better than finite-mean models. The second reason is that the sample mean of iid samples of heavy-tailed data may not converge or may even tend to infinity as the sample size increases. Therefore, it is not sufficient to conclude that infinite-mean models are unrealistic by the finiteness of the sample mean. Indeed, models with infinite moments are not “improper" as emphasized by Mandelbrot, (1997), and they have been extensively used in the financial and economic literature (see Mandelbrot, (1997) and Cont (2001)).
This paper focuses on establishing some stochastic dominance relations for infinite-mean models, applicable in several contexts. For two random variables $X$ and $Y$, $X$ is said to be stochastically smaller than (or stochastically dominated by) $Y$, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}.$ The order $\leq_{\rm st}$ is arguably the strongest form of commonly used stochastic orders; see Levy (2016) for their applications in decision making and Müller and Stoyan, (2002) and Shaked and Shanthikumar (2007) for the mathematics. Let $X$ be a Pareto random variable with infinite mean and $X_{1},\dots,X_{n}$ be iid copies of $X$. For a nonnegative vector $\left(\theta_{1},\dots,\theta_{n}\right)$ with $\sum_{i=1}^{n}\theta_{i}=1$, Chen et al. (2024a) showed that
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (1)
The inequality is strict if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive. The case of two Pareto random variables with tail parameter 1/2 has been studied in Example 7 of Embrechts et al., (2002); see Section 3 for the precise definition of the Pareto distribution. A closely related result of Ibragimov, (2005) implies that (1) also holds if $X_{1},\dots,X_{n}$ are iid positive one-sided stable random variables with infinite mean.
Inequality (1) provides very strong implications in decision making as it surprisingly holds in the strongest form of risk comparison. Intuitively, if $X_{1},\dots,X_{n}$ are treated as losses in a portfolio selection problem, any agent who prefers less loss will choose to take one of $X_{1},\dots,X_{n}$ instead of allocating their risk exposure over different losses. Other applications of (1) include optimal bundling problems (Ibragimov and Walden, (2010)), risk management (Chen et al. (2024a)), and risk sharing (Chen et al. (2024b)). It is worth noting that (1) cannot hold if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive and $X_{1},\dots,X_{n}$ have finite mean (see Proposition 2 of Chen et al. (2024a)).
Theorem 1 of Chen et al. (2024a) shows that (1) also holds for weakly negatively associated super-Pareto random variables $X_{1},\dots,X_{n}$. The class of super-Pareto random variables is quite broad and can be obtained by applying increasing and convex transforms to a Pareto random variable with tail parameter $1$. Important examples of super-Pareto distributions include the Pareto, generalized Pareto, Burr, paralogistic, and log-logistic distributions, all with infinite mean.
Given the applicability of (1), this paper aims to further generalize the inequality in two aspects: the marginal distribution and the dependence structure of $(X_{1},\dots,X_{n})$. We will first show that (1) holds for a rather large class of distributions, which we call super heavy-tailed (Theorem 1). The class of super heavy-tailed distributions includes super-Pareto distributions and has several nice properties (Propositions 2 and 3), leading to a very rich class of models; Section 3 gives many examples of super heavy-tailed distributions.
With a slight modification of the proof of Theorem 1, we show in Theorem 2 that (1) holds for negatively lower orthant dependent (Block et al., (1982)) super heavy-tailed random variables $X_{1},\dots,X_{n}$. Negative lower orthant dependence has been widely used in statistics and operations research and it relates to several popular notions of negative dependence in the literature such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)). As the class of super heavy-tailed distributions includes super-Pareto distributions and negative lower orthant dependence is implied by weak negative association, Theorem 2 implies Theorem 1 (i) of Chen et al. (2024a). Remarkably, while Theorem 2 is more general, it is shown by a much more concise proof than Chen et al. (2024a).
We proceed to study (1) given non-identically distributed random variables $X_{1},\dots,X_{n}$. Since $X_{1},\dots,X_{n}$ do not follow the same distribution, the choice of $X$ becomes unclear. A possible choice is to let $X$ follow the generalized mean of the distributions of $X_{1},\dots,X_{n}$; note that the generalized mean of distribution functions is a distribution function. A special case is the arithmetic mean, which leads to the commonly used distribution mixture models. Considering a sufficiently large class of distributions, Theorem 3 shows that (1) holds if the distribution of $X$ is the generalized mean with non-negative power of the distributions of $X_{1},\dots,X_{n}$. To our best knowledge, Theorem 3 is the first attempt to establish a non-trivial version of (1) for non-identically distributed random variables.
The rest of the paper is organized as follows. In Section 2, we present some first observations on (1). We introduce the class of super heavy-tailed distributions and establish our main result (Theorem 1) in Section 3. Section 4 presents several generalizations of Theorem 1 to negative dependence and non-identical marginal distributions. Several applications of (1) are briefly discussed in Section 5. Section 6 concludes the paper. The appendix contains the proof of Proposition 5.
1.1 Notation, conventions and definitions
In this section, we collect some notation and conventions used throughout the rest of the paper and remind the reader of some well-known definitions.
A function $f$ on $(0,\infty)$ is said to be sub-additive if $f(x+y)\leq f(x)+f(y)$ for any $x,y>0$. If the inequality is strict, we say $f$ is strictly sub-additive. For a random variable $X\sim F$, denote by $\operatorname*{ess\text{-}inf}X$ ($\operatorname*{ess\text{-}inf}F$) and $\operatorname*{ess\text{-}sup}X$ ($\operatorname*{ess\text{-}sup}F$) its essential infimum and essential supremum. Denote by $\Delta_{n}$ the standard simplex, that is, $\Delta_{n}=\{\bar{\theta}\in[0,1]^{n}:\sum_{i=1}^{n}\theta_{i}=1\}$, where we use notation $\bar{\theta}$ for a vector $(\theta_{1},\dots,\theta_{n})$. We will also use $[n]$ to denote the set of indices $1,\dots,n$.
Definition 1.
We say that a random variable $X$ is stochastically smaller than (or stochastically dominated by) a random variable $Y$, if
$\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)\quad\text{for all }x\in\mathbb{R}.$
We use notation $X\leq_{\rm st}Y$ for this property and refer to this order as stochastic order; see Shaked and Shanthikumar (2007) for properties of stochastic order. We write $X<_{\rm st}Y$ if $\mathbb{P}(X\leq x)>\mathbb{P}(Y\leq x)$ for all $x>\operatorname*{ess\text{-}inf}X$.
2 Some observations on the stochastic dominance
Throughout the paper, we work with random variables which are almost surely non-negative.
The main focus of the paper is on studying random variables $X$ such that
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n},$ (2)
for all $\bar{\theta}\in\Delta_{n}$, where $X_{1},\dots,X_{n}$ are iid copies of $X$. We will also say that a distribution $F$ satisfies property (2) if a random variable $X\sim F$ satisfies it.
Since (2) holds if a constant is added to $X$, we will, without loss of generality, only consider random variables with essential infimum 00. To avoid trivialities, we also assume throughout that $n\geq 2$ and $\theta_{i}>0$ for all $i\in[n]$. We will also be interested in distributions, and random variables, for which property (2) holds with a strict inequality. Let us start by formulating and providing some straightforward observations of (2).
Proposition 1.
Assume that random variables $X$ and $Y$ satisfy property (2) and are independent. Then the following statements hold.
  1. (i)
    $\mathbb{E}(X)=\infty$.
  2. (ii)
    A random variable $aX+b$ with $a\geq 0$ and $b\in\mathbb{R}$ satisfies (2).
  3. (iii)
    Random variables $\max\{X,c\}$ and $\max\{X,Y\}$ satisfy (2), with $c\geq 0$.
  4. (iv)
    A random variable $g(X)$ with a convex non-decreasing function $g$ satisfies (2). In addition, if $X$ satisfies (2) with a strict inequality, $g$ is convex and strictly increasing, then $g(X)$ also satisfies (2) with a strict inequality.
Proof.
  1. (i)
    Proposition 2 of Chen et al. (2024a) shows that if $X,X_{1},\dots,X_{n}$ are identically distributed with finite mean, (2) holds if and only if $X_{1}=\dots=X_{n}$ almost surely. The desired result follows.
  2. (ii)
    The proof is straightforward and is omitted.
  3. (iii)
    We will prove only the stronger property for the maximum of two random variables. Let $X_{1},\dots,X_{n}$ be iid copies of $X$, $Y_{1},\dots,Y_{n}$ be iid copies of $Y$, and $\{X_{i}\}_{i\in[n]}$ and $\{Y_{i}\}_{i\in[n]}$ be independent. For $x\in\mathbb{R}$ and $\bar{\theta}\in\Delta_{n}$, we have
    $\mathbb{P}(\max\{X,Y\}\leq x)$ $=\mathbb{P}(X\leq x)\mathbb{P}(Y\leq x)\geq\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}Y_{i}\leq x\right)$
    $=\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x,\sum_{i=1}^{n}\theta_{i}Y_{i}\leq x\right)=\mathbb{P}\left(\max\left\{\sum_{i=1}^{n}\theta_{i}X_{i},\sum_{i=1}^{n}\theta_{i}Y_{i}\right\}\leq x\right)$
    $\geq\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}\max\{X_{i},Y_{i}\}\leq x\right).$
  4. (iv)
    Since $g$ is convex and non-decreasing, $g(X)\leq_{\rm st}g(\sum_{i=1}^{n}\theta_{i}X_{i})\leq\sum_{i=1}^{n}\theta_{i}g(X_{i})$, where the first inequality holds as stochastic order is preserved under monotone transforms and the second inequality is to be understood in the almost sure (and therefore also stochastic) sense and is due to convexity of $g$. ∎
Remark 1.
Properties (ii)-(iv) above demonstrate that, even if one knows only several random variables satisfying (2), it is possible to construct many more. Of special interest is property (iii), which does not require any specific distributional properties of $X$ and $Y$ apart from property (2).
3 Super heavy-tailed distributions and stochastic dominance
In this section, we introduce a class of distributions we refer to as super heavy-tailed. We then prove that all such distributions satisfy property (2). We also explore the properties of these distributions and demonstrate that many well-known distributions with infinite mean belong to this class. Along with the results of Proposition 1, this shows that the class of distributions satisfying property (2) is large.
As has already been noted, we can, without loss of generality, consider random variables whose essential infimum is zero. For a random variable $X\sim F$ with $\operatorname*{ess\text{-}inf}X=0$, we have $F(x)>0$ for all $x>0$.
Definition 2.
Let $F$ be a distribution function with $\operatorname*{ess\text{-}inf}F=0$ and $h_{F}(x)=-\log F\left(1/x\right)$ for $x\in(0,\infty)$. If $h_{F}$ is (strictly) sub-additive, we say $F$ is (strictly) super heavy-tailed. We also say a random variable $X\sim F$ is (strictly) super heavy-tailed if $F$ is (strictly) super heavy-tailed.
Theorem 1.
If a random variable $X$ is super heavy-tailed, then it satisfies property (2). If $X$ is strictly super heavy-tailed, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
Proof.
Let $X\sim F$ and $\bar{\theta}\in\Delta_{n}$. We have, for all $x>0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)=\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
$=\exp\left(-\sum_{i=1}^{n}h_{F}\left(\frac{\theta_{i}}{x}\right)\right)\leq\exp\left(-h_{F}\left(\sum_{i=1}^{n}\frac{\theta_{i}}{x}\right)\right)=\exp\left(-h_{F}\left(\frac{1}{x}\right)\right)=F(x).$
The strictness statement is straightforward. The proof is complete. ∎
An immediate consequence of Theorem 1 and Proposition 1 (i) is that super heavy-tailed distributions have infinite mean.
3.1 Properties of super heavy-tailed distributions
In this section, we study some properties of super heavy-tailed random variables. In particular, these properties will be used to demonstrate that many well-known infinite-mean random variables are super heavy-tailed and, thanks to further sufficient conditions for belonging to this class, one can construct many more super heavy-tailed random variables.
Proposition 2.
Let $X\sim F$ be a super heavy-tailed random variable. The following statements hold.
  1. (i)
    $F$ is continuous on $[0,\infty)$.
  2. (ii)
    $F^{\beta}$, $\beta>0$, is super heavy-tailed.
  3. (iii)
    If, in addition, a random variable $Y\sim G$ is super heavy-tailed and independent of $X$, then $\max\{X,Y\}$ is also super heavy-tailed. In terms of distribution functions, if $F$ and $G$ are both super heavy-tailed, then so is $FG$.
  4. (iv)
    For a non-decreasing, convex, and non-constant function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ with $f(0)=0$, $f(X)$ is super heavy-tailed.
Proof.
  1. (i)
    As $h_{F}$ is sub-additive and increasing, and $\lim_{x\downarrow 0}h_{F}(x)=0$, $h_{F}$ is continuous on $(0,\infty)$, and so is $F$ (see Remark 1 of Matkowski and Świątkowski, (1993)). The desired result is due to the right-continuity of $F$.
  2. (ii)
    Proof of (ii) is straightforward and thus omitted.
  3. (iii)
    This is also straightforward.
  4. (iv)
    For $y\geq 0$, let $f^{-1+}(y)=\inf\{x\geq 0:f(x)>y\}$ be the right-continuous generalized inverse of $f$ with the convention that $\inf\emptyset=\infty$. As $f$ is increasing, convex, and non-constant with $f(0)=0$, $f^{-1+}$ is strictly increasing and concave and $f^{-1+}(0)\geq 0$. Therefore, by concavity of $f^{-1+}$ and $f^{-1+}(0)\geq 0$, it is clear that $f^{-1+}(tx)\geq tf^{-1+}(x)$ for any $x>0$ and $t\in(0,1]$. For any $a,b>0$,
    $f^{-1+}\left(\frac{ab}{a+b}\right)\left(f^{-1+}\left(a\right)+f^{-1+}\left(b\right)\right)$ $\geq\frac{a}{a+b}f^{-1+}\left(b\right)f^{-1+}\left(a\right)+\frac{b}{a+b}f^{-1+}\left(a\right)f^{-1+}\left(b\right)$
    $=f^{-1+}\left(a\right)f^{-1+}\left(b\right).$
    Hence, we have
    $\left(f^{-1+}\left(\frac{ab}{a+b}\right)\right)^{-1}\leq\left(f^{-1+}\left(a\right)\right)^{-1}+\left(f^{-1+}\left(b\right)\right)^{-1}.$ (3)
    Denote by $F$ and $G$ the distribution functions of $X$ and $f(X)$, respectively. Then $G(x)=\mathbb{P}(f(X)\leq x)=\mathbb{P}(X\leq f^{-1+}(x))=F(f^{-1+}(x))$ for $x\geq 0$. By letting $g(x)=1/f^{-1+}(1/x)$ for $x>0$, we write $h_{G}=h_{F}\circ g$. By inequality (3), for any $x,y>0$,
    $g(x+y)=\left(f^{-1+}\left(\frac{1/xy}{1/x+1/y}\right)\right)^{-1}\leq\left(f^{-1+}\left(\frac{1}{x}\right)\right)^{-1}+\left(f^{-1+}\left(\frac{1}{y}\right)\right)^{-1}=g(x)+g(y).$
    Therefore, $g$ is sub-additive. As $h_{F}$ is sub-additive and non-decreasing, it is clear that $h_{G}=h_{F}\circ g$ is sub-additive and we have the desired result. ∎
Proposition 3.
Let $\bar{\theta}\in\Delta_{n}$. If distribution functions $F_{1},\dots,F_{n}$ are super heavy-tailed and $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{n}$, then $\sum_{i=1}^{n}\theta_{i}F_{i}$ is also super heavy-tailed.
Proof.
Let $G=\sum_{i=1}^{n}\theta_{i}F_{i}$. It suffices to show
$G\left(\frac{xy}{x+y}\right)\geq G(x)G(y)~{}~{}~{}\mbox{for all $x,y>0$.}$ (4)
For $n=2$, as $F_{1}$ and $F_{2}$ are super heavy-tailed,
$G\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$ $=\theta_{1}F_{1}\left(\frac{xy}{x+y}\right)+\theta_{2}F_{2}\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$
$\geq\theta_{1}F_{1}\left(x\right)F_{1}\left(y\right)+\theta_{2}F_{2}\left(x\right)F_{2}\left(y\right)-G\left(x\right)G\left(y\right)$
$=\theta_{1}F_{1}\left(x\right)F_{1}\left(y\right)+\theta_{2}F_{2}\left(x\right)F_{2}\left(y\right)$
$~{}~{}~{}-(\theta_{1}F_{1}(x)+\theta_{2}F_{2}(x))(\theta_{1}F_{1}(y)+\theta_{2}F_{2}(y))$
$=\theta_{1}\theta_{2}(F_{1}\left(x\right)-F_{2}(x))(F_{1}\left(y\right)-F_{2}(y))\geq 0.$
Hence, (4) holds for $n=2$. Next, assume that (4) holds for $n=k-1$ where $k>3$ is an integer. Let $a=\sum_{i=1}^{k-1}\theta_{i}F_{i}(x)$, $b=\sum_{i=1}^{k-1}\theta_{i}F_{i}(y)$, $c=a/(F_{n}(x)(1-\theta_{n}))$, and $d=b/(F_{n}(y)(1-\theta_{n}))$. For $n=k$,
$G\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$ $=\sum_{i=1}^{k}\theta_{i}F_{i}\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$
$=\sum_{i=1}^{k-1}\theta_{i}F_{i}\left(\frac{xy}{x+y}\right)+\theta_{n}F_{n}\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$
$=(1-\theta_{n})\sum_{i=1}^{k-1}\frac{\theta_{i}}{1-\theta_{n}}F_{i}\left(\frac{xy}{x+y}\right)+\theta_{n}F_{n}\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$
$\geq(1-\theta_{n})\left(\sum_{i=1}^{k-1}\frac{\theta_{i}}{1-\theta_{n}}F_{i}\left(x\right)\right)\left(\sum_{i=1}^{k-1}\frac{\theta_{i}}{1-\theta_{n}}F_{i}\left(y\right)\right)$
$~{}~{}~{}+\theta_{n}F_{n}\left(\frac{xy}{x+y}\right)-G\left(x\right)G\left(y\right)$
$\geq\frac{ab}{1-\theta_{n}}+\theta_{n}F_{n}(x)F_{n}(y)-(a+\theta_{n}F_{n}(x))(b+\theta_{n}F_{n}(y))$
$=\frac{ab\theta_{n}}{1-\theta_{n}}+(\theta_{n}-\theta_{n}^{2})F_{n}(x)F_{n}(y)-a\theta_{n}F_{n}(y)-b\theta_{n}F_{n}(x)$
$=\theta_{n}(1-\theta_{n})F_{n}(x)F_{n}(y)\left(cd+1-c-d\right).$
As $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{k}$, $F_{k}\leq\sum_{i=1}^{k-1}\theta_{i}/(1-\theta_{n})F_{i}$. Thus $c,d\geq 1$ and $cd+1-c-d\geq 0$. The proof is complete by induction. ∎
Remark 2.
Let $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ be some function. It is well-known that if $f(x)/x$ is decreasing, $f$ is sub-additive (Theorem 7.2.4 of Hille and Phillips, (1996)). Another sufficient condition for $f$ being sub-additive is that $f$ is concave and $\lim_{x\downarrow 0}f(x)\geq 0$ (Theorem 7.2.5 of Hille and Phillips, (1996)). Hence, if $h_{F}(x)/x$ is decreasing or $h_{F}$ is concave, $F$ is super heavy-tailed.
3.2 Examples of super heavy-tailed distributions
In this section, we demonstrate that many well-known infinite-mean distributions belong to the super heavy-tailed class.
Example 1 (Pareto distribution).
For $\alpha>0$, the Pareto distribution, denoted by Pareto$(\alpha)$, is defined as
$\mathbb{P}(X\leq x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
If $\alpha\leq 1$, the Pareto distribution is super heavy-tailed. This can be seen by noting that any Pareto$(\alpha)$ random variable $X$ can be written as $X=f(Z)$ where $Z\sim{\rm Pareto}(1)$ and $f(x)=(x+1)^{1/\alpha}-1$ for $x\geq 0$. It is easy to show that Pareto$(1)$ is super heavy-tailed. As $f$ is increasing and convex if $\alpha\leq 1$, the result follows from Proposition 2 (iv).
Example 2 (Fréchet distribution).
For $\alpha>0$, the Fréchet distribution is defined as
$\mathbb{P}(X\leq x)=\exp(-x^{-\alpha}),~{}~{}x>0.$
If $\alpha\leq 1$, the Fréchet distribution is super heavy-tailed. To see this, let $F$ denote the distribution function of $X$. We have $h_{F}(x)=x^{\alpha}$ for $x>0$. Then, if $\alpha\leq 1$, for $x,y>0$,
$\frac{h_{F}(x)+h_{F}(y)}{h_{F}(x+y)}=\left(\frac{x}{x+y}\right)^{\alpha}+\left(1-\frac{x}{x+y}\right)^{\alpha}\geq 1.$
Thus $h_{F}$ is sub-additive and the Fréchet distribution with $\alpha\leq 1$ is super heavy-tailed.
Example 3 (Generalized Pareto distribution).
The generalized Pareto distribution with parameters $\xi\in\mathbb{R}$ and $\beta>0$ is defined as
$\mathbb{P}(X\leq x)=\begin{cases}1-\left(1+\xi\frac{x}{\beta}\right)^{-1/\xi}&x\in D(\xi,\beta)\\ e^{-x/\beta}&x\geq 0,\end{cases}$
where $D(\xi,\beta)=[0,\infty)$ if $\xi>0$ and $D(\xi,\beta)=[0,-\beta/\xi)$ if $\xi<0$. By the Pickands-Balkema-de Haan Theorem (Balkema and de Haan,, 1974; Pickands,, 1975), the generalized Pareto distributions are the only possible non-degenerate limiting distributions of the excess of random variables beyond a high threshold. If $\xi\geq 1$, then the generalized Pareto distribution is super heavy-tailed. This is by Proposition 2 (iv); that is, the generalized Pareto random variables with $\xi\geq 1$ can be obtained from location-scale transforms of Pareto$(1/\xi)$ random variables.
Example 4 (Super-Pareto distribution).
Introduced by Chen et al. (2024a), a super-Pareto distribution is the distribution of a random variable $f(X+1)$ where $X\sim\rm{Pareto}(1)$ and $f$ is some non-decreasing, convex, and non-constant function. By Proposition 2 (iv), a super-Pareto distribution (after some location-scale transform) is super heavy-tailed.
Example 5 (Burr distribution).
For $\alpha,\tau>0$, the Burr distribution is defined as
$\mathbb{P}(X\leq x)=1-\left(\frac{1}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0.$ (5)
Let $Y\sim\rm{Pareto}(\alpha)$. Then $Y^{1/\tau}$ follows a Burr distribution. If $\alpha,\tau\leq 1$, the Burr distribution is super-Pareto and hence super heavy-tailed. Special cases of Burr distributions are the paralogistic ($\alpha=\tau$) and the log-logistic ($\alpha=1$) distributions; see Kleiber and Kotz, (2003) and Klugman et al. (2012).
Example 6 (Inverse Burr distribution).
Suppose that $Y$ follows the Burr distribution (5). Then $X=1/Y$ follows the inverse Burr distribution
$\mathbb{P}(X\leq x)=\left(\frac{x^{\tau}}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0,$
where $\alpha,\tau>0$. Let $F$ denote the inverse Burr distribution function. If $\tau\leq 1$, it is easy to check that the second derivative of $h_{F}$ is always negative, and thus $h_{F}$ is sub-additive (see Remark 2). Hence $X$ is super heavy-tailed if $\tau\leq 1$. Note that the super heavy-tailed property may not always be preserved under the inverse transformation of super heavy-tailed random variables. For instance, if $Z$ follows a Fréchet distribution without finite mean, then $1/Z$ follows a Weibull distribution whose mean is always finite and it cannot be super heavy-tailed.
Example 7 (Log-Pareto distribution).
If $Y\sim\mathrm{Pareto}(\alpha)$, $\alpha>0$, then $X=\exp(Y)-1$ has a log-Pareto distribution (see p. 39 in Arnold, (2015)), with distribution function
$\mathbb{P}(X\leq x)=1-\frac{1}{(\log(x+1)+1)^{\alpha}},~{}~{}x>0.$
If $\alpha\in(0,1]$, by Proposition 2 (iv), $X$ is super heavy-tailed.
Example 8 (Stoppa distribution).
For $\alpha>0$ and $\beta>0$, a (location-shifted) Stoppa distribution can be defined as
$\mathbb{P}(X\leq x)=\left(1-\frac{1}{(x+1)^{\alpha}}\right)^{\beta},~{}~{}x>0.$
Since a Stoppa distribution is a power transform of a Pareto distribution, by Proposition 2 (ii), if $\alpha\leq 1$, Stoppa distribution is super heavy-tailed. Power transforms have also been used to generalize Burr distributions (see p. 211 of Kleiber and Kotz, (2003)).
Example 9 (Mixture of distributions).
Suppose that $F_{1},\dots,F_{n}$ are Pareto distributions with possibly different tail parameters $0<\alpha_{1},\dots,\alpha_{n}\leq 1$. As $F_{1},\dots,F_{n}$ are comparable in stochastic order, by Proposition 3, mixtures of $F_{1},\dots,F_{n}$ are super heavy-tailed. Similarly, mixtures of super heavy-tailed Stoppa distributions are also super heavy-tailed.
Example 10 (Excess of loss).
For a super heavy-tailed random variable $X$, by Proposition 2 (iv), $(X-m)_{+}$ is also super heavy-tailed where $m>0$. Here, $(X-m)_{+}$ can be interpreted as the loss assumed by a (re)insurer who issues excess-of-loss (re)insurance contracts with retention level $m$.
4 Weighted sums of non-iid random variables
So far we have focused on property (2) where $X_{1},\dots,X_{n}$ are independent copies of $X$. In this section, we explore what happens for $X_{1},\dots,X_{n}$ which are negatively dependent or do not necessarily have the same marginal distribution.
4.1 Negatively dependent random variables
The notion of negative dependence below will be used to establish the main result of this section.
Definition 3 (Block et al., (1982)).
Random variables $X_{1},\dots,X_{n}$ are negatively lower orthant dependent (NLOD) if for all $x_{1},\dots,x_{n}\in\mathbb{R}$, $\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\leq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i})$.
Negative lower orthant dependence is implied by several popular notions of negative dependence in the literature, such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)), negative orthant dependence (Block et al., (1982)), and negative regression dependence (Lehmann (1966) and Block et al., (1985)). It is also implied by a recently introduced notion called weak negative association (Chen et al. (2024a)). See Chen et al. (2024a) for more details about the implication relations of these notions of negative dependence. Examples of NLOD random variables include those following normal distributions with non-positive correlations (see Joag-Dev and Proschan (1983)).
A straightforward observation allows one to prove a version of Theorem 1 for NLOD super heavy-tailed random variables.
Theorem 2.
If a random variable $X$ is super heavy-tailed and random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, then for $\bar{\theta}\in\Delta_{n}$,
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (6)
If $X$ is strictly super heavy-tailed, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
Proof.
We provide a proof here for completeness but it is a minor variation of the proof of Theorem 1. Let $X\sim F$. We have, for all $x>0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
$=\exp\left(-\sum_{i=1}^{n}h_{F}\left(\frac{\theta_{i}}{x}\right)\right)\leq\exp\left(-h_{F}\left(\sum_{i=1}^{n}\frac{\theta_{i}}{x}\right)\right)=\exp\left(-h_{F}\left(\frac{1}{x}\right)\right)=F(x).$
The proof is complete. ∎
Inequality (6) was established in Theorem 1 (i) of Chen et al. (2024a) for super-Pareto random variables $X_{1},\dots,X_{n}$ being weakly negatively associated. As weakly negatively associated super-Pareto random variables (after location-scale transforms) are super heavy-tailed and NLOD, Theorem 2 generalizes Theorem 1 (i) of Chen et al. (2024a). We present below a corollary that generalizes Theorem 1 (ii) of Chen et al. (2024a).
Corollary 1.
Suppose that a random variable $X$ is super heavy-tailed, random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, and $\xi_{1},\dots,\xi_{n}$ are any positive random variables independent of $X,X_{1},\dots,X_{n}$.
  1. (i)
    We have $\sum_{i=1}^{n}\xi_{i}X\leq_{\rm st}\sum_{i=1}^{n}\xi_{i}X_{i}$.
  2. (ii)
    If, in addition, $\mathbb{P}(cX>t)\geq c\mathbb{P}(X>t)$ for all $c\in(0,1]$ and $t>0$, and $\sum_{i=1}^{n}\xi_{i}\leq 1$, then for all $x\geq 0$,
    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}(X>x).$ (7)
Proof.
  1. (i)
    By Theorem 2 and the independence between $\xi_{1},\dots,\xi_{n}$ and $X,X_{1},\dots,X_{n}$, for $x\in\mathbb{R}$,
    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X>x\right).$
    This implies (i).
  2. (ii)
    With the additional assumptions of (ii), we can again write
    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right].$
    Thanks to (6), the RHS is greater than or equal to
    $\mathbb{E}\left[\mathbb{P}\left(\left(\sum_{i=1}^{n}\xi_{i}\right)X>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}\left(X>x\right).$
    The proof is complete. ∎
Note that for Corollary 1 to hold, no dependence assumptions are needed between $\xi_{1},\dots,\xi_{n}$. We explain below how Theorem 1 (ii) of Chen et al. (2024a) can be implied by (7). For $\bar{\theta}\in\Delta_{n}$, let $A_{1},\dots,A_{n}$ be any events independent of $(X_{1},\dots,X_{n})$ and event $A$ be independent of $X$ satisfying $\mathbb{P}(A)=\sum_{i=1}^{n}\theta_{i}\mathbb{P}(A_{i})$. If $X_{1},\dots,X_{n}$ are financial losses, $A_{1},\dots,A_{n}$ can be interpreted as the triggering events for these losses. Let $\xi_{i}=\theta_{i}\mathds{1}_{A_{i}}$. By (7), for $x\geq 0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\theta_{i}\mathds{1}_{A_{i}}\right)\mathbb{P}(X>x)=\mathbb{P}(A)\mathbb{P}(X>x)=\mathbb{P}(X\mathds{1}_{A}>x).$
Equivalently, we can write the above inequality as $X\mathds{1}_{A}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}$. This result was shown in Theorem 1 (ii) of Chen et al. (2024a) for weakly negatively associated and regular super-Pareto random variables, which satisfy the assumptions of Corollary 1 (ii); we refer to Chen et al. (2024a) for the definition of regular super-Pareto random variables. Thus we have the implication.
The following corollary shows that the stochastic dominance (6) holds strictly for Pareto distributions with infinite mean.
Corollary 2.
Suppose that $X,X_{1},\dots,X_{n}\sim{\rm Pareto}(\alpha)$ with $\alpha\leq 1$ and $X_{1},\dots,X_{n}$ are NLOD. Then for $\bar{\theta}\in\Delta_{n}$,
$X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}.$ (8)
Proof.
Let $F={\rm Pareto}(\alpha)$ with $\alpha\leq 1$. By Theorem 2, it suffices to show $h_{F}$ is strictly sub-additive. Equivalently, we need to show $D(x,y):=F(1/(x+y))-F(1/x)F(1/y)>0$ for $x,y>0$. We have $D(x,y)=xy/((x+1)(y+1)(x+y+1))>0$. As Pareto random variables with parameter $\alpha<1$ can be obtained from Pareto random variables with $\alpha=1$ by strictly increasing and convex transforms (see Example 1), the desired result follows from Proposition 1 (iv). ∎
The strict inequality (8) plays an important role in many optimization problems (see, e.g., risk exchange in Chen et al. (2024b) and optimal bundling in Ibragimov and Walden, (2010)). We can see from the proof that as many random variables can be obtained by applying strictly increasing and convex transforms to Pareto$(1)$ random variables, (8) holds for a large class of random variables (see Section 3.2). A similar proof shows that Fréchet random variables with a tail parameter strictly less than 1 also satisfy (8).
Remark 3.
Suppose $X_{ij}$ for all $i\in[m]$ and $j\in[n]$ are super heavy-tailed random variables such that $X_{i1},\dots,X_{in}$ are independent for each $i\in[m]$, and $X_{1j},\dots,X_{mj}$ are NLOD and identically distributed for each $j\in[n]$. Let $Y_{i}=\sum_{j=1}^{n}X_{ij}$ for all $i\in[m]$. As stochastic order is closed under convolution (e.g., Theorem 1.A.3 of Shaked and Shanthikumar (2007)), for $(\theta_{1},\dots,\theta_{m})\in\Delta_{m}$, $Y_{1}\leq_{\rm st}\sum_{i=1}^{m}\theta_{i}Y_{i}.$
Remark 4.
One may expect positive dependence to make larger values of the sum in (6) more likely and thus the sum more likely to stochastically dominate a single random variable. We believe that this intuition does not hold due to the very heavy tails of the random variables under consideration. It is known, for instance, that very large values of the sum of iid random variables with heavy tails are usually caused by a single random variable taking a large value, while other random variables are moderate. If random variables are positively dependent and some of them do not take large values, it makes others more likely to take moderate values too, hence positive dependence hinders large values. The situation is of course reversed for negative dependence. However, (6) can still hold for Pareto random variables with infinite mean that are positively dependent via some specific Clayton copula (see Chen et al., 2024c ).
4.2 Non-identical distributions
In the previous sections, stochastic dominance relations are obtained for random variables with the same marginal distribution. We now look at the case when the random variables are not necessarily identically distributed. Given non-identically distributed random variables $X_{1},\dots,X_{n}$ and any $\bar{\theta}\in\Delta_{n}$, the question is to study for which random variable $X$ the following property holds
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (9)
To study this problem, we consider the class of super-Fréchet distributions defined below.
Definition 4.
A random variable $X$ with $\operatorname*{ess\text{-}inf}X=0$ is said to be super-Fréchet (or has a super-Fréchet distribution) if the function $g:x\mapsto 1/(-\log\mathbb{P}(X\leq x))$ is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g(x)=0$.
Super-Fréchet random variables can be obtained by applying strictly increasing and convex transforms to Fréchet random variables with tail parameter 1, and therefore are super heavy-tailed. As convex transforms make the tail of random variables heavier, super-Fréchet distributions are more heavy-tailed than Fréchet distribution with tail parameter 1, and thus the name. Fréchet distributions with infinite mean, as well as many other distributions in the following example, are super-Fréchet.
Example 11.
Pareto, Burr, paralogistic, and log-logistic random variables, all with infinite mean, are super-Fréchet distributions. Since all these random variables can be obtained by applying strictly increasing and convex transforms to Pareto random variables with tail parameter $1$ (see Example 4), it suffices to show that a Pareto random variable with tail parameter $1$ is super-Fréchet. Write the Pareto distribution with tail parameter 1 as
$F(x)=1-\frac{1}{x+1}=\exp\left(-\frac{1}{g(x)}\right),~{}~{}~{}x>0,$
where $g(x)=1/\log(1+1/x)$. It is clear that $g$ is strictly increasing and $\lim_{x\downarrow 0}g(x)=0$. We show $g$ is concave on $(0,\infty)$. We have
$g^{\prime\prime}(x)=\frac{2-(1+2x)\log(1+1/x)}{x^{2}(1+x)^{2}\log^{3}(1+1/x)}.$
Let $r(x)=\log\left(1/x+1\right)-2/(1+2x)$, $x>0$. It is easy to verify that $r$ is strictly decreasing on $(0,\infty)$ and $r(x)$ goes to 00 as $x$ goes to infinity. Thus $r(x)>0$ and $g^{\prime\prime}(x)<0$ for $x\in(0,\infty)$.
We will assume $X_{1},\dots,X_{n}$ in (9) are super-Fréchet. Since $X_{1},\dots,X_{n}$ may not have the same distribution, how to choose the distribution of $X$ is not clear. A perhaps natural candidate is the generalized mean of the distributions of $X_{1},\dots,X_{n}$. For $r\in\mathbb{R}\setminus\{0\}$, $n\in\mathbb{N}$, and $\mathbf{w}=(w_{1},\dots,w_{n})\in\Delta_{n}$, the generalized $r$-mean function is defined as
$M^{\mathbf{w}}_{r}(u_{1},\dots,u_{n})=\left(w_{1}u_{1}^{r}+\dots+w_{n}u_{n}^{r}\right)^{1/r},~{}~{}~{}~{}~{}(u_{1},\dots,u_{n})\in(0,\infty)^{n}.$
The generalized 00-mean function is the weighted geometric mean, that is, $M^{\mathbf{w}}_{0}(u_{1},\dots,u_{n})=\prod_{i=1}^{n}u_{i}^{w_{i}},$ which is also the limit of $M^{\mathbf{w}}_{r}$ as $r\to 0$. A generalized mean of distribution functions is a distribution function. In particular, if $r=1$, it leads to a distribution mixture model.
Theorem 3.
Suppose that $X_{1},\dots,X_{n}$ are super-Fréchet, NLOD, and $X_{i}\sim F_{i}$, $i\in[n]$. For $\bar{\theta}\in\Delta_{n}$ and $0\leq t\leq s$,
$X^{*}_{s}\leq_{\rm st}X^{*}_{t}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i},$
where $X^{*}_{r}\sim M^{\bar{\theta}}_{r}(F_{1},\dots,F_{n})$, $r\geq 0$.
Proof.
Let $g_{i}(x)=1/(-\log F_{i}(x))$, $x>0$, for all $i\in[n]$. As $g_{i}$, $i\in[n]$, is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g_{i}(x)=0$, $g_{i}(x)\geq\theta g_{i}(x/\theta)$ for all $x>0$ and $\theta\in(0,1)$. Then, for $\theta\in(0,1)$,
$F_{i}\left(\frac{x}{\theta}\right)=\exp\left(-g_{i}\left(\frac{x}{\theta}\right)^{-1}\right)\leq\exp\left(-\theta g_{i}(x)^{-1}\right)=F_{i}(x)^{\theta}.$ (10)
As $X_{1},\dots,X_{n}$ are NLOD, by (10), for any $x>0$, $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$, and $r\geq 0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F_{i}\left(\frac{x}{\theta_{i}}\right)\leq\prod_{i=1}^{n}F_{i}\left(x\right)^{\theta_{i}}$
$=M^{\bar{\theta}}_{0}(F_{1}(x),\dots,F_{n}(x))\leq M^{\bar{\theta}}_{r}(F_{1}(x),\dots,F_{n}(x))=\mathbb{P}(X^{*}_{r}\leq x).$
The last inequality is because the generalized mean function is monotone in $r$; that is, given any $\mathbf{w}\in\Delta_{n}$, $M^{\mathbf{w}}_{r}\leq M^{\mathbf{w}}_{s}$ for $r\leq s$ (Theorem 16 of Hardy et al. (1934)). ∎
5 Applications
This section is devoted to some applications of the stochastic dominance properties studied in this paper.
5.1 Portfolio diversification
Since the introduction of the modern portfolio theory (Markowitz (1952)), it is well known that diversification can reduce risks of decision makers in finance and insurance sectors, especially when losses have finite means (Samuelson, (1967)). However, diversification can also be detrimental in the presence of extremely heavy-tailed losses, as warned by Fama and Miller, (1972), Embrechts et al., (2002), and Ibragimov, (2005). We illustrate below the effects of diversification on super heavy-tailed losses.
Suppose that an agent needs to allocate their exposure $\theta\in\mathbb{R}_{+}$ over NLOD and identically distributed super heavy-tailed losses $X_{1},\dots,X_{n}$ to minimize the risk. They use a risk measure to evaluate their risk; a risk measure is a functional $\rho:\mathcal{X}_{\rho}\rightarrow\mathbb{R}$ where $\mathcal{X}_{\rho}$ is a set of random losses. The two regulatory risk measures in insurance and finance are Value-at-Risk (VaR) and Expected Shortfall (ES). For a random variable $X\sim F$ and $p\in(0,1)$, VaR is defined as the left quantile
$\mathrm{VaR}_{p}(X)=F^{-1}(p)=\inf\{t\in\mathbb{R}:F(t)\geq p\},$
and ES is defined as
$\mathrm{ES}_{p}(X)=\frac{1}{1-p}\int_{p}^{1}\mathrm{VaR}_{u}(X)\mathrm{d}u.$
Note that for losses with infinite mean, some risk measures such as ES, may give infinite value and thus are not useful for our analysis. For the rest of our discussion, we assume $\mathcal{X}_{\rho}$ contains the convex cone generated by $X_{1},\dots,X_{n}$. For two random losses $X$ and $Y$, an agent equipped with a risk measure $\rho$ prefers $X$ over $Y$ if $\rho(X)\leq\rho(Y)$. All commonly used risk measures, including VaR and ES, are monotone, that is, $\rho(X)\leq\rho(Y)$ if $X\leq_{\rm st}Y$; see Föllmer and Schied (2016) for a detailed discussion on risk measures. The result below directly follows from Theorem 2.
Proposition 4.
Suppose that $X_{1},\dots,X_{n}$ are identically distributed, NLOD, and super heavy-tailed. For $(\theta_{1},\dots,\theta_{n})\in\mathbb{R}_{+}^{n}$ such that $\sum_{i=1}^{n}\theta_{i}=\theta$ and a monotone risk measure $\rho:\mathcal{X}_{\rho}\rightarrow\mathbb{R}$, we have
$\rho(\theta X_{1})\leq\rho\left(\sum_{i=1}^{n}\theta_{i}X_{i}\right).$ (11)
Inequality (11) suggests a perhaps surprising implication: The agent should not diversify over super heavy-tailed losses if they want to minimize their risk, which does not depend on their risk preference. This is in sharp contrast to the case of iid finite-mean losses, for which diversification is beneficial.
If $\rho$ is taken to be VaR, then (11) implies the super-additivity of VaR: For all $p\in(0,1)$,
$\mathrm{VaR}_{p}(\theta_{1}X_{1})+\dots+\mathrm{VaR}_{p}(\theta_{n}X_{n})\leq\mathrm{VaR}_{p}(\theta_{1}X_{1}+\dots+\theta_{n}X_{n}).$ (12)
Although (12) is not new, the previous results (e.g., Embrechts et al. (2009)) usually hold in some asymptotic sense (i.e., $p$ is close to 1) rather than for all $p\in(0,1)$.
5.2 An optimal bundling problem
Bundling is a market strategy to sell goods as collections. We complement and generalize the results of the Vickrey auction model studied by Palfrey, (1983) and Ibragimov and Walden, (2010). Suppose that there is a single seller who provides $m$ goods to $n$ buyers in the market. Let $2^{M}$ be the power set of $[m]$. A bundling decision $\beta=\{B_{1},\dots,B_{l}\}$ is a partition of $[m]$ such that
  1. (a)
    $B_{s}\neq\emptyset$ for any $s\in[l]$,
  2. (b)
    $B_{s}\cap B_{t}=\emptyset$, for any distinct $s,t\in[l]$
  3. (c)
    $\bigcup_{s=1}^{l}B_{s}=[m]$.
For $j\in[n]$, the $j$th buyer’s valuations for goods are represented by $\mathbf{X}_{j}=(X_{1j},\dots,X_{mj})$ where $X_{ij}$ is their valuation for good $i\in[m]$. Assume that $X_{1j},\dots,X_{mj}$ are identically distributed and NLOD super heavy-tailed random variables. The valuations of buyers are independent of each other, i.e., $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ are independent. Each buyer knows their valuations of the goods but not those of the other buyers; the only available information is the distributions of the other buyers’ valuations. Assume that each buyer’s valuation for bundles is additive. That is, for a bundle $B\in 2^{M}$, its valuation of buyer $j$ is $v_{j}(B)=\sum_{i\in B}X_{ij}$ and for bundling decision $\beta=(B_{1},\dots,B_{l})$, its valuation of buyer $j$ is $v_{j}(\beta)=\sum_{s=1}^{l}v_{j}(B_{s})$.
Let the seller distribute its goods via Vickrey auctions. In this situation, all buyers will simultaneously submit sealed bids for bundles of goods. Whoever submits the highest bid will win the goods and pay the seller the second-highest bid. For buyer $j\in[n]$, let $\mathbf{x}_{j}=(x_{1j},\dots,x_{mj})\in\mathbb{R}_{+}^{m}$ be their valuations for goods. Note that the valuations of the other buyers are unknown to buyer $j$, and thus are random variables. That is, to buyer $j$, $v_{j}(B)=\sum_{i\in B}x_{ij}$ and $v_{s}(B)=\sum_{i\in B}X_{is}$ for any $s\in[n]/\{j\}$ and bundle $B\in 2^{M}$. Then the expected surplus of buyer $j$ with valuations $\mathbf{x}_{j}$ for bundle $B\in 2^{M}$ is
$\mathbb{E}(S_{j}(B,\mathbf{x}_{j}))=$ $\mathbb{P}\left(\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)\mathbb{E}\left(v_{j}(B)-\max_{s\in[n]/\{j\}}v_{s}(B)|\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)$
The expected surplus of buyer $j$ with valuations $\mathbf{x}_{j}$ for bundling decision $\beta=(B_{1},\dots,B_{l})$ is $\mathbb{E}(S_{j}(\beta,\mathbf{x}_{j}))=\sum_{s=1}^{l}\mathbb{E}(S_{j}(B_{s},\mathbf{x}_{j}))$. We say buyer $j$ with valuations $\mathbf{x}_{j}$ prefers a bundling decision $\beta_{1}$ to a bundling decision $\beta_{2}$ if $\mathbb{E}(S_{j}(\beta_{1},\mathbf{x}_{j}))\geq\mathbb{E}(S_{j}(\beta_{2},\mathbf{x}_{j}))$. Buyers unanimously prefer $\beta_{1}$ to $\beta_{2}$ if, for almost all of their valuations, they prefer $\beta_{1}$ to $\beta_{2}$.
Proposition 5.
Suppose that the valuations of goods are identically distributed, NLOD, and super heavy-tailed. Then the buyers unanimously prefer unbundled sales $\{\{1\},\{2\},\dots,\{m\}\}$.
That is, the buyers prefer separate auctions if their valuations are super heavy-tailed. In contrast, the buyers never prefer separate auctions if the valuations are bounded and there are more than two buyers, as shown by Palfrey, (1983). If the valuations are iid positive stable random variables with infinite mean, the statement in Proposition 5 also holds (Theorem 4.1 of Ibragimov and Walden, (2010)). Thus, our result generalizes that of Ibragimov and Walden, (2010) in both the distributions and dependence of valuations for goods; Banciu and Ødegaard, (2016) studied the pricing of a bundle of products with dependent valuations. Proposition 5 can be shown following the proof of Theorem 4.1 of Ibragimov and Walden, (2010). For completeness, we provide a proof in the Appendix.
5.3 A single-period newsvendor problem
As one of the classical problems in inventory management, the single-period newsvendor problem is to find the optimal level of inventory such that the expected profit of a store is maximized. We consider the most basic single-period newsvendor problem for illustration; recent advances in the literature can be found in Qin et al., (2011).
At the beginning of the period, the customer demand for a single product $D$ is assumed to be a random variable with density $f$ and distribution $F$ and it will be realized at the end of the period. A store needs to determine an inventory level $q$ of the products with fixed cost $v$ per unit. At the end of the period, if $D\leq q$, then $q-D$ units which are left over are salvaged by the store for revenue of $g<v$ per unit. If $D>q$, then $D-q$ units are “lost" sales cost the store $b$ per unit. Let the price of the product $p$ be fixed. Assume $g=b=0$ for simplicity. Then the expected profit of the store is
$\mathbb{E}(\Pi(D,q))=$ $\int_{0}^{q}(px-vq)f(x)\mathrm{d}x+\int_{q}^{\infty}(pq-vq)f(x)\mathrm{d}x=p\mathbb{E}(\min(D,q))-vq.$ (13)
It is well known that the optimal inventory level which maximizes (13) is (e.g., Qin et al., (2011))
$F^{-1}\left(\frac{p-v}{p}\right)=\operatorname*{arg\,max}_{q\in\mathbb{R}}\{\mathbb{E}(\Pi(D,q))\},$
where $(p-v)/p$ is called the margin ratio.
Now assume that the store provides a single product to $2$ customers with demands $X_{1}\sim F_{1}$ and $X_{2}\sim F_{2}$. The store can either (1) stock the products at 2 locations to serve the customers separately or (2) stock products together in one location (inventory pooling). The optimal inventory levels of the two strategies are thus (1) $F_{1}^{-1}((p-v)/p)+F_{2}^{-1}((p-v)/p)$ and (2) $H^{-1}((p-v)/p)$, where $H$ is the distribution of $X_{1}+X_{2}$. It has been observed in the literature that inventory pooling can reduce the inventory level in many model assumptions; see Eppen, (1979). However, inventory pooling can sometimes increase the inventory level (i.e., $F_{1}^{-1}((p-v)/p)+F_{2}^{-1}((p-v)/p)\leq H^{-1}((p-v)/p)$), and this is called “inventory anomaly" in Yang and Schrage, (2009).
Proposition 6.
Suppose that the demands of the two customers are identically distributed, NLOD, and super heavy-tailed. Then inventory pooling always increases the optimal inventory level.
Proof.
This is a direct result from Theorem 2 and (1.A.12) of Shaked and Shanthikumar (2007). ∎
A similar result was shown in Aydın et al., (2012) assuming that the demands follow a regularly varying distribution and the tail probability of their joint distribution is negligible in some sense. Their result requires additional constraints on the margin ratio whereas the above proposition holds for any margin ratios between 0 and 1. Another close observation is discussed following Theorem 1 of Bimpikis and Markakis, (2016) for iid demands which follow stable distributions with infinite mean.
6 Conclusion
In this paper, we provide some sufficient conditions for property (2) to hold. One can see that the property, while very strong, holds for a remarkably large class of distributions. We have also shown that it remains valid for random variables which are negatively dependent or non-identically distributed.
We conclude with some open questions which are topics of our ongoing research. First, we are interested in understanding how close our sufficient conditions for (2) are to the optimal ones, i.e., we would like to understand what conditions are necessary for (2).
Second, the definition of super heavy-tailed random variables seems to suggest that it is the distribution of $1/X$ that is of importance. We currently lack an intuitive explanation of this.
Finally, property (2) raises the possibility that, for some random variables $X_{1},\dots,X_{n}$ and two vectors $\bar{\eta},\bar{\gamma}\in\mathbb{R}^{n}_{+}$,
$\eta_{1}X_{1}+\dots+\eta_{n}X_{n}\leq_{\rm st}\gamma_{1}X_{1}+\dots+\gamma_{n}X_{n},$ (14)
where $\bar{\gamma}$ is smaller than $\bar{\eta}$ in majorization order; that is, $\sum^{n}_{i=1}\gamma_{i}=\sum^{n}_{i=1}\eta_{i}$ and $\sum^{k}_{i=1}\gamma_{(i)}\geq\sum^{k}_{i=1}\eta_{(i)}\,$ for $k\in[n-1]$ where $\gamma_{(i)}$ and $\eta_{(i)}$ represent the $i$th smallest order statistics of $\bar{\gamma}$ and $\bar{\eta}$. Clearly, (14) implies (2). Inequality (14) has been shown in Chen et al., 2024c for iid Pareto random variables with infinite mean. A part of our ongoing research is to show that (14) in fact holds for a much larger class of distributions. Note that the methods used in the current paper do not appear to be useful to address (14) as we rely on the comparison of a sum with each of the summands. A more subtle approach to sums is required.
Appendices
Appendix A Proof of Proposition 5
Suppose $n\geq 2$. Let $\mathbf{x}_{j}=(x_{1j},\dots,x_{mj})\in\mathbb{R}_{+}^{m}$ be the valuations of buyer $j\in[n]$ and $v_{j}(B)=\sum_{i\in B}x_{ij}$, $B\in 2^{M}$. Let $(X_{1s},\dots,X_{ms})$ be the valuations of buyer $s\in[n]/\{j\}$ and $X_{is}\sim F$, $i\in[m]$. For $B\in 2^{M}$, denote by $H$ the distribution function of $v_{s}(B)=\sum_{i\in[B]}X_{is}$. Then by independence of $v_{s}(B)$, $s\in[n]/\{j\}$,
$\mathbb{P}\left(\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)=H\left(v_{j}(B)\right)^{n-1}.$
Using integration by parts, we get
$\mathbb{E}\left(\max_{s\in[n]/\{j\}}v_{s}(B)|\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)$ $=\int_{0}^{v_{j}(B)}x\mathrm{d}\left(\left(\frac{H\left(x\right)}{H\left(v_{j}(B)\right)}\right)^{n-1}\right)$
$=v_{j}(B)-\int_{0}^{v_{j}(B)}\left(\frac{H\left(x\right)}{H\left(v_{j}(B)\right)}\right)^{n-1}\mathrm{d}x.$
By Theorem 1, $H(kx)\leq F(x)$, where $k=\mathrm{card}(B)$. Hence the expected surplus for buyer $j$ with valuations $\mathbf{x}_{j}$ is
$\mathbb{E}(S_{j}(B,\mathbf{x}_{j}))=\int_{0}^{v_{j}(B)}H\left(x\right)^{n-1}\mathrm{d}x=k\int_{0}^{v_{j}(B)/k}H\left(kx\right)^{n-1}\mathrm{d}x,\leq k\int_{0}^{v_{j}(B)/k}F(x)^{n-1}\mathrm{d}x.$
As $F(x)^{n-1}$ is increasing for $x\geq 0$, $\int_{0}^{y}F(x)^{n-1}\mathrm{d}x$ is a convex function of $y$. By Theorem 3.C.1 of Marshall et al. (2011), $F(y_{1},\dots,y_{k})=\sum_{t=1}^{k}\int_{0}^{y_{t}}F(x)^{n-1}\mathrm{d}x$, $(y_{1},\dots,y_{k})\in\mathbb{R}_{+}^{k}$, is Schur-convex, which implies $F(\sum_{t=1}^{k}y_{t}/k,\dots,\sum_{t=1}^{k}y_{t}/k)\leq F(y_{1},\dots,y_{k})$. Hence,
$k\int_{0}^{v_{j}(B)/k}F(x)^{n-1}\mathrm{d}x$ $=\sum_{t\in B}\int_{0}^{\sum_{t\in B}x_{tj}/k}F(x)^{n-1}\mathrm{d}x$
$\leq\sum_{t\in B}\int_{0}^{x_{tj}}F(x)^{n-1}\mathrm{d}x=\sum_{t\in B}\mathbb{E}(S_{j}(\{t\},\mathbf{x}_{j})).$
Since $\mathbf{x}_{j}$ is arbitrary and the expected surplus for bundling decisions is additive, the proof is done.
Acknowledgements
The authors thank Qihe Tang for the helpful discussions and comments that led to this paper.

### v3

@@
-Stochastic dominance for super heavy-tailed random variables
+Risk aggregation and stochastic dominance for a class of heavy-tailed distributions
 Abstract
-We introduce a class of super heavy-tailed distributions and establish the inequality that any weighted average of independent and identically distributed super heavy-tailed random variables stochastically dominates one such random variable.
-We show that many commonly used extremely heavy-tailed (i.e., infinite-mean) distributions, such as the Pareto, Fréchet, and Burr distributions, belong to the class of super heavy-tailed distributions.
-The established stochastic dominance relation is further generalized to allow negatively dependent or non-identically distributed random variables.
-In particular, the weighted average of non-identically distributed random variables stochastically dominates their distribution mixtures.
-Applications of these results in portfolio diversification, goods bundling, and inventory management are discussed.
-Remarkably, in the presence of super heavy-tailedness, the results that hold for finite-mean models in these applications are flipped.
+We introduce a new class of heavy-tailed distributions for which any weighted average of independent and identically distributed random variables is larger than one such random variable in (usual) stochastic order.
+We show that many commonly used extremely heavy-tailed (i.e., infinite-mean) distributions, such as the Pareto, Fréchet, and Burr distributions, belong to this class.
+The established stochastic dominance relation can be further generalized to allow negatively dependent or non-identically distributed random variables.
+In particular, the weighted average of non-identically distributed random variables dominates their distribution mixtures in stochastic order.
 Keywords: heavy-tailed distributions; stochastic order; negative dependence; infinite mean.
@@
 Distributions with infinite mean are ubiquitous in the realm of banking and insurance, and they are particularly useful in modeling catastrophic losses (Ibragimov et al.
-(2009)), operational losses (Moscadelli, (2004)), costs of cyber risk events (Eling and Wirfs, (2019)), and financial returns from technology innovations (Silverberg and Verspagen, (2007)); see also Chen et al.
-(2024b) for a list of empirical examples of distributions with infinite mean.
+(2009)), operational losses (Moscadelli, (2004)), costs of cyber risk events (Eling and Wirfs, (2019)), and financial returns from technology innovations (Silverberg and Verspagen, (2007)); see also Chen and Wang (2025) for a list of empirical examples of distributions with infinite mean.
 As the world is arguably finite (e.g., any loss is bounded by the total wealth in the world), why should we use models with infinite mean as mathematical tools?
-There are perhaps two reasons.
-The first reason is that, after careful statistical analysis, infinite-mean models often fit extremely heavy-tailed datasets better than finite-mean models.
-The second reason is that the sample mean of iid samples of heavy-tailed data may not converge or may even tend to infinity as the sample size increases.
+The main reason is that infinite-mean models often fit extremely heavy-tailed datasets better than finite-mean models.
+Moreover, the sample mean of iid samples of heavy-tailed data may not converge or may even tend to infinity as the sample size increases.
 Therefore, it is not sufficient to conclude that infinite-mean models are unrealistic by the finiteness of the sample mean.
 Indeed, models with infinite moments are not “improper" as emphasized by Mandelbrot, (1997), and they have been extensively used in the financial and economic literature (see Mandelbrot, (1997) and Cont (2001)).
-This paper focuses on establishing some stochastic dominance relations for infinite-mean models, applicable in several contexts.
-For two random variables $X$ and $Y$, $X$ is said to be stochastically smaller than (or stochastically dominated by) $Y$, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}.$ The order $\leq_{\rm st}$ is arguably the strongest form of commonly used stochastic orders; see Levy (2016) for their applications in decision making and Müller and Stoyan, (2002) and Shaked and Shanthikumar (2007) for the mathematics.
-Let $X$ be a Pareto random variable with infinite mean and $X_{1},\dots,X_{n}$ be iid copies of $X$.
-For a nonnegative vector $\left(\theta_{1},\dots,\theta_{n}\right)$ with $\sum_{i=1}^{n}\theta_{i}=1$, Chen et al.
-(2024a) showed that
+This paper focuses on establishing some stochastic dominance relations for infinite-mean models.
+For two random variables $X$ and $Y$, $X$ is said to be smaller than $Y$ in stochastic order, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}$; see Müller and Stoyan, (2002) and Shaked and Shanthikumar (2007) for extensive accounts of properties of stochastic dominance.
+Let $X$ be a positive one-sided stable random variable with infinite mean and $X_{1},\dots,X_{n}$ be iid copies of $X$.
+For a nonnegative vector $\left(\theta_{1},\dots,\theta_{n}\right)$ with $\sum_{i=1}^{n}\theta_{i}=1$, Ibragimov, (2005) showed that
 $X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (1)
-The inequality is strict if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive.
-The case of two Pareto random variables with tail parameter 1/2 has been studied in Example 7 of Embrechts et al., (2002); see Section 3 for the precise definition of the Pareto distribution.
-A closely related result of Ibragimov, (2005) implies that (1) also holds if $X_{1},\dots,X_{n}$ are iid positive one-sided stable random variables with infinite mean.
+Recently, Arab et al., (2024), Chen et al.
+(2024a), and Müller, (2024) have shown that inequality (1) holds for more general classes of distributions.
+The case of two Pareto random variables with tail parameter 1/2 was studied in Example 7 of Embrechts et al., (2002); see Section 3 for the precise definition of the Pareto distribution.
 Inequality (1) provides very strong implications in decision making as it surprisingly holds in the strongest form of risk comparison.
-Intuitively, if $X_{1},\dots,X_{n}$ are treated as losses in a portfolio selection problem, any agent who prefers less loss will choose to take one of $X_{1},\dots,X_{n}$ instead of allocating their risk exposure over different losses.
-Other applications of (1) include optimal bundling problems (Ibragimov and Walden, (2010)), risk management (Chen et al.
-(2024a)), and risk sharing (Chen et al.
+If $X_{1},\dots,X_{n}$ are treated as losses in a portfolio selection problem, any agent who prefers less loss will choose to take one of $X_{1},\dots,X_{n}$ instead of allocating their risk exposure over different losses.
+This observation is counterintuitive, contrasting with the common belief that diversification reduces risk.
+Other applications of (1) include optimal bundling problems (Ibragimov and Walden, (2010)) and risk sharing (Chen et al.
 (2024b)).
-It is worth noting that (1) cannot hold if at least two components of $\left(\theta_{1},\dots,\theta_{n}\right)$ are positive and $X_{1},\dots,X_{n}$ have finite mean (see Proposition 2 of Chen et al.
-(2024a)).
-Theorem 1 of Chen et al.
-(2024a) shows that (1) also holds for weakly negatively associated super-Pareto random variables $X_{1},\dots,X_{n}$.
+In this paper, we will study (1) where $X_{1},\dots,X_{n}$ are possibly negatively dependent, a case not considered in Ibragimov, (2005), Arab et al., (2024), and Müller, (2024).
+Chen et al.
+(2024a) have shown that (1) also holds for weakly negatively associated super-Pareto random variables $X_{1},\dots,X_{n}$.
 The class of super-Pareto random variables is quite broad and can be obtained by applying increasing and convex transforms to a Pareto random variable with tail parameter $1$.
-Important examples of super-Pareto distributions include the Pareto, generalized Pareto, Burr, paralogistic, and log-logistic distributions, all with infinite mean.
-Given the applicability of (1), this paper aims to further generalize the inequality in two aspects: the marginal distribution and the dependence structure of $(X_{1},\dots,X_{n})$.
-We will first show that (1) holds for a rather large class of distributions, which we call super heavy-tailed (Theorem 1).
-The class of super heavy-tailed distributions includes super-Pareto distributions and has several nice properties (Propositions 2 and 3), leading to a very rich class of models; Section 3 gives many examples of super heavy-tailed distributions.
-With a slight modification of the proof of Theorem 1, we show in Theorem 2 that (1) holds for negatively lower orthant dependent (Block et al., (1982)) super heavy-tailed random variables $X_{1},\dots,X_{n}$.
-Negative lower orthant dependence has been widely used in statistics and operations research and it relates to several popular notions of negative dependence in the literature such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)).
-As the class of super heavy-tailed distributions includes super-Pareto distributions and negative lower orthant dependence is implied by weak negative association, Theorem 2 implies Theorem 1 (i) of Chen et al.
-(2024a).
-Remarkably, while Theorem 2 is more general, it is shown by a much more concise proof than Chen et al.
-(2024a).
-We proceed to study (1) given non-identically distributed random variables $X_{1},\dots,X_{n}$.
+Examples of super-Pareto distributions include the Pareto, generalized Pareto, Burr, paralogistic, and log-logistic distributions, all with infinite mean.
+This paper aims to further generalize the result of Chen et al.
+(2024a) in two aspects: the marginal distribution and the dependence structure of $(X_{1},\dots,X_{n})$.
+In Section 3, we first introduce a new class of distributions, which has several nice properties (Propositions 2 and 3) and includes the class of super-Pareto distributions as a special case.
+Within this class of distributions, we show in Theorem 1 that (1) holds for identically distributed random variables $X_{1},\dots,X_{n}$ that are negatively lower orthant dependent (Block et al., (1982)).
+It is well known that the behavior of the sum of extremely heavy-tailed random variables is dominated by the maximum of the summands (see Embrechts et al.
+(1997)).
+Therefore, a possible reason why (1) is preserved when transitioning from independence to negative dependence is because under negative dependence, random variables that take small to moderate values will push the other random variables to take large values with a larger probability, leading to a stochastically larger $\sum_{i=1}^{n}\theta_{i}X_{i}$.
+The situation is different for positively dependent random variables; see Remark 5.
+As negative lower orthant dependence is more general than weak negative association, Theorem 1 (i) of Chen et al.
+(2024a) is implied by Theorem 1.
+Remarkably, while Theorem 1 is more general, it is shown by a much more concise proof.
+In Section 4, we proceed to study (1) given non-identically distributed random variables $X_{1},\dots,X_{n}$.
 Since $X_{1},\dots,X_{n}$ do not follow the same distribution, the choice of $X$ becomes unclear.
-A possible choice is to let $X$ follow the generalized mean of the distributions of $X_{1},\dots,X_{n}$; note that the generalized mean of distribution functions is a distribution function.
+A possible choice is to let $X$ follow the generalized mean of the distributions of $X_{1},\dots,X_{n}$.
 A special case is the arithmetic mean, which leads to the commonly used distribution mixture models.
-Considering a sufficiently large class of distributions, Theorem 3 shows that (1) holds if the distribution of $X$ is the generalized mean with non-negative power of the distributions of $X_{1},\dots,X_{n}$.
+Considering a rather large class of distributions, Theorem 2 shows that (1) holds if the distribution of $X$ is the generalized mean with non-negative power of the distributions of $X_{1},\dots,X_{n}$.
 To our best knowledge, Theorem 2 is the first attempt to establish a non-trivial version of (1) for non-identically distributed random variables.
@@
 In Section 2, we present some first observations on (1).
-We introduce the class of super heavy-tailed distributions and establish our main result (Theorem 1) in Section 3.
-Section 4 presents several generalizations of Theorem 1 to negative dependence and non-identical marginal distributions.
-Several applications of (1) are briefly discussed in Section 5.
+Sections 3 and 4 present the main results.
+Section 5 compares our results with the literature.
 Section 6 concludes the paper.
-The appendix contains the proof of Proposition 5.
+The appendix contains the proofs of Propositions 2 and 3 as well as some examples in the new class of distributions.
 1.1 Notation, conventions and definitions
 In this section, we collect some notation and conventions used throughout the rest of the paper and remind the reader of some well-known definitions.
-A function $f$ on $(0,\infty)$ is said to be sub-additive if $f(x+y)\leq f(x)+f(y)$ for any $x,y>0$.
-If the inequality is strict, we say $f$ is strictly sub-additive.
+A function $f$ on $(0,\infty)$ is said to be subadditive if $f(x+y)\leq f(x)+f(y)$ for any $x,y>0$.
+If the inequality is strict, we say $f$ is strictly subadditive.
 For a random variable $X\sim F$, denote by $\operatorname*{ess\text{-}inf}X$ ($\operatorname*{ess\text{-}inf}F$) and $\operatorname*{ess\text{-}sup}X$ ($\operatorname*{ess\text{-}sup}F$) its essential infimum and essential supremum.
 Denote by $\Delta_{n}$ the standard simplex, that is, $\Delta_{n}=\{\bar{\theta}\in[0,1]^{n}:\sum_{i=1}^{n}\theta_{i}=1\}$, where we use notation $\bar{\theta}$ for a vector $(\theta_{1},\dots,\theta_{n})$.
+Let $\Delta_{n}^{+}=\Delta_{n}\cap(0,1)^{n}$.
 We will also use $[n]$ to denote the set of indices $1,\dots,n$.
+For a distribution function $F$, its generalized inverse is defined as
+$F^{-1}(p)=\inf\{t\in\mathbb{R}:F(t)\geq p\},~{}p\in(0,1).$
 Definition 1.
-We say that a random variable $X$ is stochastically smaller than (or stochastically dominated by) a random variable $Y$, if
-$\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)\quad\text{for all }x\in\mathbb{R}.$
-We use notation $X\leq_{\rm st}Y$ for this property and refer to this order as stochastic order; see Shaked and Shanthikumar (2007) for properties of stochastic order.
-We write $X<_{\rm st}Y$ if $\mathbb{P}(X\leq x)>\mathbb{P}(Y\leq x)$ for all $x>\operatorname*{ess\text{-}inf}X$.
+We say that a random variable $X$ is smaller than a random variable $Y$ in stochastic order, denoted by $X\leq_{\rm st}Y$, if $\mathbb{P}(X\leq x)\geq\mathbb{P}(Y\leq x)$ for all $x\in\mathbb{R}$.
+For random variables $X$ and $Y$ with support $[c,\infty)$ where $c\in\mathbb{R}$, we write $X<_{\rm st}Y$ if $\mathbb{P}(X\leq x)>\mathbb{P}(Y\leq x)$ for all $x>c$.
 2 Some observations on the stochastic dominance
@@
 The main focus of the paper is on studying random variables $X$ such that
-$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n},$ (2)
-for all $\bar{\theta}\in\Delta_{n}$, where $X_{1},\dots,X_{n}$ are iid copies of $X$. We will also say that a distribution $F$ satisfies property (2) if a random variable $X\sim F$ satisfies it.
-Since (2) holds if a constant is added to $X$, we will, without loss of generality, only consider random variables with essential infimum 00.
-To avoid trivialities, we also assume throughout that $n\geq 2$ and $\theta_{i}>0$ for all $i\in[n]$.
-We will also be interested in distributions, and random variables, for which property (2) holds with a strict inequality.
-Let us start by formulating and providing some straightforward observations of (2).
+$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}\mbox{~{}for all $\bar{\theta}\in\Delta_{n}$,}$ (SD)
+where $X_{1},\dots,X_{n}$ are independent or negatively dependent with the marginal laws equal to $X$ (see Section 3.2 for the precise definition of negative dependence).
+We will also say that a distribution $F$ satisfies property (SD) if a random variable $X\sim F$ satisfies it.
+If some of $\theta_{1},\dots,\theta_{n}$ are 0, we can simply reduce the dimension of our problem.
+Therefore, for most of our results, we will assume $\bar{\theta}\in\Delta_{n}^{+}$.
+Since (SD) holds if a constant is added to $X$, we will, without loss of generality, only consider random variables with essential infimum 00.
+We will also be interested in distributions, and random variables, for which property (SD) holds with a strict inequality.
+Let us start by formulating and providing some straightforward observations of (SD).
 Proposition 1.
-Assume that random variables $X$ and $Y$ satisfy property (2) and are independent. Then the following statements hold.
+Assume that random variables $X$ and $Y$ satisfy property (SD) and are independent. Then the following statements hold.
   1. (i)
-    $\mathbb{E}(X)=\infty$.
+    $\mathbb{E}(X)=\infty$ or $X$ is a constant.
   2. (ii)
-    A random variable $aX+b$ with $a\geq 0$ and $b\in\mathbb{R}$ satisfies (2).
+    A random variable $aX+b$ with $a\geq 0$ and $b\in\mathbb{R}$ satisfies (SD).
   3. (iii)
-    Random variables $\max\{X,c\}$ and $\max\{X,Y\}$ satisfy (2), with $c\geq 0$.
+    Random variables $\max\{X,c\}$ and $\max\{X,Y\}$ satisfy (SD), with $c\geq 0$.
   4. (iv)
-    A random variable $g(X)$ with a convex non-decreasing function $g$ satisfies (2).
-In addition, if $X$ satisfies (2) with a strict inequality, $g$ is convex and strictly increasing, then $g(X)$ also satisfies (2) with a strict inequality.
+    A random variable $g(X)$ with a convex non-decreasing function $g$ satisfies (SD).
+In addition, if $X$ satisfies (SD) with a strict inequality, $g$ is convex and strictly increasing, then $g(X)$ also satisfies (SD) with a strict inequality.
 Proof.
   1. (i)
-    Proposition 2 of Chen et al.
-(2024a) shows that if $X,X_{1},\dots,X_{n}$ are identically distributed with finite mean, (2) holds if and only if $X_{1}=\dots=X_{n}$ almost surely.
-The desired result follows.
+    This is implied by Proposition 2 of Chen et al. (2024a).
   2. (ii)
@@
     We will prove only the stronger property for the maximum of two random variables.
-Let $X_{1},\dots,X_{n}$ be iid copies of $X$, $Y_{1},\dots,Y_{n}$ be iid copies of $Y$, and $\{X_{i}\}_{i\in[n]}$ and $\{Y_{i}\}_{i\in[n]}$ be independent.
+Let $X_{1},\dots,X_{n}$ follow the distribution of $X$, $Y_{1},\dots,Y_{n}$ follow the distribution of $Y$, and $\{X_{i}\}_{i\in[n]}$ and $\{Y_{i}\}_{i\in[n]}$ be independent.
 For $x\in\mathbb{R}$ and $\bar{\theta}\in\Delta_{n}$, we have
@@
   4. (iv)
-    Since $g$ is convex and non-decreasing, $g(X)\leq_{\rm st}g(\sum_{i=1}^{n}\theta_{i}X_{i})\leq\sum_{i=1}^{n}\theta_{i}g(X_{i})$, where the first inequality holds as stochastic order is preserved under monotone transforms and the second inequality is to be understood in the almost sure (and therefore also stochastic) sense and is due to convexity of $g$.
+    Since $g$ is convex and non-decreasing, $g(X)\leq_{\rm st}g(\sum_{i=1}^{n}\theta_{i}X_{i})\leq\sum_{i=1}^{n}\theta_{i}g(X_{i})$, where the first inequality holds as stochastic order is preserved under non-decreasing transforms and the second inequality is to be understood in the almost sure (and therefore also stochastic) sense and is due to convexity of $g$.
 ∎
-Remark 1.
-Properties (ii)-(iv) above demonstrate that, even if one knows only several random variables satisfying (2), it is possible to construct many more.
-Of special interest is property (iii), which does not require any specific distributional properties of $X$ and $Y$ apart from property (2).
-3 Super heavy-tailed distributions and stochastic dominance
-In this section, we introduce a class of distributions we refer to as super heavy-tailed.
-We then prove that all such distributions satisfy property (2).
-We also explore the properties of these distributions and demonstrate that many well-known distributions with infinite mean belong to this class.
-Along with the results of Proposition 1, this shows that the class of distributions satisfying property (2) is large.
+Properties (ii)-(iv) above demonstrate that, even if one knows only several random variables satisfying (SD), it is possible to construct many more.
+Of special interest is property (iii), which does not require any specific distributional properties of $X$ and $Y$ apart from property (SD).
+3 A class of heavy-tailed distributions and stochastic dominance
+In this section, we introduce a new class of heavy-tailed distributions.
+We explore several properties of this class and demonstrate that it contains many well-known distributions with infinite mean.
+We then prove that all distributions in this class satisfy property (SD).
+Along with the results of Proposition 1, this shows that the class of distributions satisfying property (SD) is large.
+3.1 A class of heavy-tailed distributions
 As has already been noted, we can, without loss of generality, consider random variables whose essential infimum is zero.
@@
 Definition 2.
-Let $F$ be a distribution function with $\operatorname*{ess\text{-}inf}F=0$ and $h_{F}(x)=-\log F\left(1/x\right)$ for $x\in(0,\infty)$.
-If $h_{F}$ is (strictly) sub-additive, we say $F$ is (strictly) super heavy-tailed.
-We also say a random variable $X\sim F$ is (strictly) super heavy-tailed if $F$ is (strictly) super heavy-tailed.
+Let $F$ be a distribution function with $\operatorname*{ess\text{-}inf}F=0$ and let $h_{F}(x)=-\log F(1/x)$ for $x\in(0,\infty)$.
+We say that $F$ belongs to $\mathcal{H}$, denoted by $F\in\mathcal{H}$, if $h_{F}$ is subadditive.
+We write $F\in\mathcal{H}_{s}$ if $h_{F}$ is strictly subadditive.
+For $X\sim F$, we also write $X\sim\mathcal{H}$ (resp.
+$X\sim\mathcal{H}_{s}$) if $F\in\mathcal{H}$ (resp.
+$F\sim\mathcal{H}_{s}$).
+Remark 1.
+By properties of subadditive functions (e.g., Theorem 7.2.4 and Theorem 7.2.5 of Hille and Phillips, (1996)), $F\in\mathcal{H}$ if $h_{F}(x)/x$ is decreasing or $h_{F}$ is concave.
+In the case of continuous distribution $F$, $F\in\mathcal{H}$ holds if and only if the survival function of $1/X$ is log-superadditive where $X\sim F$.
+We will see later that all distributions in $\mathcal{H}$ have infinite mean and because of that we say $\mathcal{H}$ is a class of heavy-tailed distributions.
+Note that the definition of heavy-tailed distributions varies in different contexts; see, e.g., Remark 3.
+Below are some examples in class $\mathcal{H}$.
+Example 1 (Fréchet distribution).
+For $\alpha>0$, the Fréchet distribution, denoted by Fréchet$(\alpha)$, is defined as
+$F(x)=\exp(-x^{-\alpha}),~{}~{}x>0.$
+If $\alpha\leq 1$, $F$ has infinite mean. It is easy to check that $F\in\mathcal{H}$ if $\alpha\leq 1$ and $F\in\mathcal{H}_{s}$ if $\alpha<1$, since for any $x,y>0$,
+$\frac{h_{F}(x)+h_{F}(y)}{h_{F}(x+y)}=\left(\frac{x}{x+y}\right)^{\alpha}+\left(1-\frac{x}{x+y}\right)^{\alpha}\geq 1.$
+As $h_{F}$ is additive when $\alpha=1$, Fréchet$(1)$ distribution can be thought as a “boundary” of class $\mathcal{H}$.
+Example 2 (Pareto$(1)$ distribution).
+For $\alpha>0$, the Pareto distribution, denoted by Pareto$(\alpha)$, is defined as
+$F(x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
+Pareto$(\alpha)$ distributions have infinite mean if $\alpha\leq 1$.
+Taking second derivative of $h_{F}$ when $\alpha=1$, we have $h^{\prime\prime}_{F}(x)=-1/(x+1)^{2}.$ Hence $h_{F}$ is concave and ${\rm Pareto}(1)\in\mathcal{H}_{s}$.
+We can show that Pareto$(\alpha)$ with $\alpha\leq 1$, as well as many other infinite-mean distributions in Table 1 also belong to $\mathcal{H}$ either directly using the definition, or using some closure properties of $\mathcal{H}$ in Propositions 2 and 3 provided below; see Appendix for detailed derivations of examples in Table 1 and the proofs of Propositions 2 and 3.
+Distribution functions Parameters
+Fréchet distribution $F(x)=\exp(-x^{-\alpha}),~{}~{}x>0$ $\alpha\leq 1$
+Pareto distribution $F(x)=1-(x+1)^{-\alpha},~{}~{}x>0$ $\alpha\leq 1$
+Generalized Pareto distribution $F(x)=1-\left(1+\xi(x/\beta)\right)^{-1/\xi},~{}~{}x>0$ $\xi\geq 1$
+Burr distribution $F(x)=1-\left(x^{\tau}+1\right)^{-\alpha},~{}~{}x>0$ $\alpha,\tau\leq 1$
+Inverse Burr distribution $F(x)=\left(x^{\tau}/(x^{\tau}+1)\right)^{\alpha},~{}~{}x>0$ $\alpha>0$, $\tau\leq 1$
+Log-Pareto distribution $F(x)=1-(\log(x+1)+1)^{-\alpha},~{}~{}x>0.$ $\alpha\leq 1$
+Stoppa distribution $F(x)=\left(1-(x+1)^{-\alpha}\right)^{\beta},~{}~{}x>0$ $\alpha\leq 1$, $\beta>0$
+Table 1: Examples of distributions in $\mathcal{H}$
+Proposition 2.
+Let $X\sim F$ where $F\in\mathcal{H}$. The following statements hold.
+  1. (i)
+    If $F$ is strictly increasing on $[0,\infty)$ and $\mathbb{P}(X<\infty)=1$, then $F$ is continuous on $[0,\infty)$.
+  2. (ii)
+    For $\beta>0$, $F^{\beta}\in\mathcal{H}$.
+  3. (iii)
+    If, in addition, a random variable $Y\sim G$, where $G\in\mathcal{H}$, is independent of $X$, then $\max\{X,Y\}\in\mathcal{H}$.
+In terms of distribution functions, if $F,G\in\mathcal{H}$, then $FG\in\mathcal{H}$.
+  4. (iv)
+    For a non-decreasing, convex, and non-constant function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ with $f(0)=0$, $f(X)\in\mathcal{H}$.
+Proposition 3.
+Let $\bar{\theta}\in\Delta_{n}^{+}$. If distribution functions $F_{1},\dots,F_{n}\in\mathcal{H}$ and $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{n}$, then $\sum_{i=1}^{n}\theta_{i}F_{i}\in\mathcal{H}$.
+It is clear that the various transforms of distributions in Proposition 2 from our class generate many different distributions, showing that the class $\mathcal{H}$ is indeed rather large.
+Suppose that $F_{1},\dots,F_{n}$ are Pareto distributions with possibly different tail parameters $0<\alpha_{1},\dots,\alpha_{n}\leq 1$.
+As $F_{1},\dots,F_{n}$ are comparable in stochastic order, by Proposition 3, mixtures of $F_{1},\dots,F_{n}$ are in $\mathcal{H}$.
+3.2 Negative lower orthant dependence
+The notion of negative dependence below will be used to establish the main result of this section.
+Definition 3 (Block et al., (1982)).
+Random variables $X_{1},\dots,X_{n}$ are negatively lower orthant dependent (NLOD) if for all $x_{1},\dots,x_{n}\in\mathbb{R}$, $\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\leq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i})$.
+Negative lower orthant dependence includes independence as a special case.
+It is commonly used in various research areas and it is implied by other popular notions of negative dependence in the literature, such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)), negative orthant dependence (Block et al., (1982)), and negative regression dependence (Lehmann (1966) and Block et al., (1985)) see, e.g., Chi et al., (2024) for the implications of these notions.
+3.3 Main result
 Theorem 1.
-If a random variable $X$ is super heavy-tailed, then it satisfies property (2). If $X$ is strictly super heavy-tailed, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
+If a random variable $X\in\mathcal{H}$ and random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, then for $\bar{\theta}\in\Delta_{n}^{+}$,
+$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (2)
+If $X\in\mathcal{H}_{s}$, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
 Proof.
-Let $X\sim F$ and $\bar{\theta}\in\Delta_{n}$. We have, for all $x>0$,
-$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)=\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
+Let $X\sim F$ and $\bar{\theta}\in\Delta_{n}^{+}$. We have, for all $x>0$,
+$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
 $=\exp\left(-\sum_{i=1}^{n}h_{F}\left(\frac{\theta_{i}}{x}\right)\right)\leq\exp\left(-h_{F}\left(\sum_{i=1}^{n}\frac{\theta_{i}}{x}\right)\right)=\exp\left(-h_{F}\left(\frac{1}{x}\right)\right)=F(x).$
 The strictness statement is straightforward. The proof is complete. ∎
-An immediate consequence of Theorem 1 and Proposition 1 (i) is that super heavy-tailed distributions have infinite mean.
-3.1 Properties of super heavy-tailed distributions
-In this section, we study some properties of super heavy-tailed random variables.
-In particular, these properties will be used to demonstrate that many well-known infinite-mean random variables are super heavy-tailed and, thanks to further sufficient conditions for belonging to this class, one can construct many more super heavy-tailed random variables.
-Proposition 2.
-Let $X\sim F$ be a super heavy-tailed random variable. The following statements hold.
+An immediate consequence of Theorem 1 and Proposition 1 (i) is that all distributions in $\mathcal{H}$ have infinite mean.
+Remark 2 (Value-at-Risk).
+One regulatory risk measure in insurance and finance is Value-at-Risk (VaR).
+For a random variable $X\sim F$ and $p\in(0,1)$, VaR is defined as $\mathrm{VaR}_{p}(X)=F^{-1}(p)$.
+For two random variables $X$ and $Y$, it is well known that $X\leq_{\rm st}Y$ if and only if $\mathrm{VaR}_{p}(X)\leq\mathrm{VaR}_{p}(Y)$ for all $p\in(0,1)$.
+Note that VaR is comonotonic-additive; a risk measure $\rho$ is comonotonic-additive if $\rho(Y+Z)=\rho(Y)+\rho(Z)$ for comonotonic random variables $Y$ and $Z$ [^1] Random variables $Y$ and $Z$ are comonotonic if there exists a random variable $U$ and two increasing functions $f$ and $g$ such that $Y=f(U)$ and $Z=g(U)$ almost surely.
+Then we have $\mathrm{VaR}_{p}(X)=\sum_{i=1}^{n}\mathrm{VaR}_{p}(\theta_{i}X_{i})$ for identically distributed random variables $X_{1},\dots,X_{n}$.
+By Theorem 1, superadditivity of VaR holds for $\bar{\theta}\in\Delta_{n}^{+}$ and identically distributed risks $X_{1},\dots,X_{n}\in\mathcal{H}$ that are NLOD: For all $p\in(0,1)$, $\mathrm{VaR}_{p}(\theta_{1}X_{1})+\dots+\mathrm{VaR}_{p}(\theta_{n}X_{n})\leq\mathrm{VaR}_{p}(\theta_{1}X_{1}+\dots+\theta_{n}X_{n}).$ More generally, the superadditivity property holds for any comonotonic-additive risk measure that is consistent with stochastic order.
+Remark 3 (Heavy-tailed distributions).
+A distribution $F$ is said to be heavy-tailed in the sense of Falk et al., (2011) with tail parameter $\alpha>0$, if $\overline{F}(x)=L(x)x^{-\alpha}$ where $L$ is a slowly varying function, that is, $L(tx)/L(x)\rightarrow 1$ as $x\rightarrow\infty$ for all $t>0$.
+For iid heavy-tailed random variables $X_{1},X_{2},\dots\sim F$, if there exist sequences of constants $\{a_{n}\}$ and $\{b_{n}\}$ where $b_{n}>0$ such that $(\max\{X_{1},\dots,X_{n}\}-a_{n})/b_{n}$ converges to the Fréchet distribution, $F$ is said to be in the maximum domain attraction of the Fréchet distribution.
+It is known in the Extreme Value Theory (Embrechts et al.
+(1997)) that a distribution is in the maximum domain of attraction of the Fréchet distribution if and only if the distribution is heavy-tailed.
+Note that for a heavy-tailed random variable $X$ with $\alpha\leq 1$, $\mathbb{E}(|X|)=\infty$.
+An interesting property of heavy-tailed risks with infinite mean is the asymptotic superadditivity of VaR: If $X_{1},\dots,X_{n}$ are iid and heavy-tailed with tail parameter $\alpha<1$,
+$\lim_{p\rightarrow 1}\frac{\mathrm{VaR}_{p}(X_{1}+\dots+X_{n})}{\mathrm{VaR}_{p}(X_{1})+\dots+\mathrm{VaR}_{p}(X_{n})}>1.$
+See, e.g., Example 3.1 of Embrechts et al.
+(2009) for the claim above.
+Heavy-tailed risks with infinite mean are not necessarily in $\mathcal{H}$ as the condition of $\mathcal{H}$ applies over the whole range of distributions whereas heavy-tailed distributions have power-law shapes only in their tail parts.
+On the other hand, risks in $\mathcal{H}$ are not necessarily heavy-tailed in the sense of Falk et al., (2011).
+For instance, the survival distributions of log-Pareto risks are slowly varying functions.
+Distributions with slowly varying tails are called super heavy-tailed.
+Remark 4 (Convex order).
+Besides stochastic order, another popular notion of stochastic dominance to compare risks is convex order.
+For two random variables $X$ and $Y$, $X$ is said to be smaller than $Y$ in convex order, denoted by $X\leq_{\rm cx}Y$, if $\mathbb{E}(u(X))\leq\mathbb{E}(u(Y))$ for all convex functions $u$ provided that the expectations exist.
+The interpretation of $X\leq_{\rm cx}Y$ is that $Y$ is more “spread-out” than $X$.
+If $X_{1},\dots,X_{n}$ are iid and have a finite mean, by Theorem 3.A.35 of Shaked and Shanthikumar (2007), for $\bar{\theta}\in\Delta_{n}^{+}$, $\sum_{i=1}^{n}\theta_{i}X_{i}\leq_{\rm cx}X_{1}$.
+Unlike Theorem 1, this leads to a diversification benefit.
+Note that $\leq_{\rm cx}$ is not suitable for the analysis of risks with infinite mean as the expectation of any increasing convex transform of these risks is infinity.
+Remark 5 (Positive dependence).
+One may expect positive dependence to make larger values of the sum in (2) more likely and thus the sum more likely to stochastically dominate a single random variable.
+We believe that this intuition does not hold due to the very heavy tails of the random variables under consideration.
+It is known, for instance, that very large values of the sum of iid random variables with heavy tails are likely caused by a single random variable taking a large value, while other random variables are moderate.
+If random variables are positively dependent and some of them do not take large values, it makes others more likely to take moderate values too, hence positive dependence may hinder large values; see Alink et al., (2004) and Mainik and Rüschendorf, (2010) for such observations in some asymptotic senses.
+These phenomena can also be seen from the deadly risks considered by Müller, (2024): For all $i\in[n]$, $\mathbb{P}(X_{i}=0)=1-p$ and $\mathbb{P}(X_{i}=\infty)=p$ where $p>0$.
+For $\overline{\theta}\in\Delta_{n}^{+}$, it is clear that $\sum_{i=1}^{n}\theta_{i}X_{i}=\infty$ as long as one of $X_{1},\dots,X_{n}$ is $\infty$.
+If $X_{1},\dots,X_{n}$ are positively lower orthant dependent (PLOD), that is $\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\geq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i})$ for all $x_{1},\dots,x_{n}\in\mathbb{R}$, then
+$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}=\infty\right)$ $=1-\mathbb{P}\left(X_{1}=\dots=X_{n}=0\right),$
+$=1-\mathbb{P}\left(X_{1}\leq 0,\dots,X_{n}\leq 0\right)\leq 1-\prod_{i=1}^{n}\mathbb{P}(X_{i}=0).$
+Hence, $\sum_{i=1}^{n}\theta_{i}X_{i}$ is stochastically smaller when $X_{1},\dots,X_{n}$ are PLOD compared to the case when $X_{1},\dots,X_{n}$ are independent.
+The situation is reversed for NLOD random variables.
+However, (SD) still holds for PLOD risks $X_{1},\dots,X_{n}$ as
+$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}=\infty\right)=1-\mathbb{P}\left(X_{1}=\dots=X_{n}=0\right)\geq 1-\mathbb{P}(X_{1}=0)=\mathbb{P}(X_{1}=\infty).$
+In Chen et al., (2025), (SD) is shown to hold for infinite-mean Pareto random variables that are positively dependent via some specific Clayton copula.
+4 Weighted sums of non-identically distributed risks
+In the previous section, property (SD) is studied for risks with the same marginal distribution.
+We now look at the case when risks are not necessarily identically distributed.
+Given non-identically distributed random variables $X_{1},\dots,X_{n}$ and any $\bar{\theta}\in\Delta_{n}$, the question is to study for which random variable $X$ the following property holds
+$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (3)
+To study this problem, we introduce the class of super-Fréchet distributions defined below.
+Definition 4.
+A random variable $X$ is said to be super-Fréchet (or has a super-Fréchet distribution) if $X$ and $f(Y)$ have the same distribution, where $Y\sim$ Fréchet$(1)$ and $f$ is a strictly increasing and convex function with $f(0)=0$.
+As convex transforms make the tail of random variables heavier, super-Fréchet distributions are more heavy-tailed than Fréchet$(1)$ distribution, and thus the name.
+It is easy to check that a random variable $X$ with $\operatorname*{ess\text{-}inf}X=0$ is super-Fréchet if and only if the function $g:x\mapsto 1/(-\log\mathbb{P}(X\leq x))$ is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g(x)=0$.
+As Fréchet$(1)$ distribution is in $\mathcal{H}$, by Proposition 1 (iv), Super-Fréchet distributions are in $\mathcal{H}$.
+On the other hand, not all distributions in $\mathcal{H}$ are Super-Fréchet, which can be seen from the following example.
+Example 3.
+For $c>0$, define a distribution function on $(0,\infty]$:
+$F(x)=\exp(-c\lceil 1/x\rceil),\mbox{~{}~{}for~{}~{}}x\in(0,\infty),$
+and $F(\infty)=1$.
+Then $h_{F}(x)=c\lceil x\rceil$, $x\in(0,\infty)$, is subadditive, and thus $F\in\mathcal{H}$.
+However, since $F$ is not a continuous distribution, it is not super-Fréchet.
+The distributions in this example are the so-called inverse-geometric distributions, also considered in Example 2.7 of Arab et al., (2024).
+Fréchet distributions with infinite mean, as well as many other distributions in the following example, are super-Fréchet.
+Example 4.
+Pareto, Burr, paralogistic, and log-logistic random variables, all with infinite mean, are super-Fréchet distributions.
+Since all these random variables can be obtained by applying strictly increasing and convex transforms to Pareto$(1)$ random variables (see Appendix C), it suffices to show that a Pareto$(1)$ random variable is super-Fréchet.
+Write the Pareto$(1)$ distribution as $F(x)=1-1/(x+1)=\exp(-1/g(x))$, $x>0$, where $g(x)=1/\log(1+1/x)$.
+It is clear that $g$ is strictly increasing and $\lim_{x\downarrow 0}g(x)=0$.
+We show $g$ is concave on $(0,\infty)$.
+We have
+$g^{\prime\prime}(x)=\frac{2-(1+2x)\log(1+1/x)}{x^{2}(1+x)^{2}\log^{3}(1+1/x)}.$
+Let $r(x)=\log\left(1/x+1\right)-2/(1+2x)$, $x>0$.
+It is easy to verify that $r$ is strictly decreasing on $(0,\infty)$ and $r(x)$ goes to 00 as $x$ goes to infinity.
+Thus $r(x)>0$ and $g^{\prime\prime}(x)<0$ for $x\in(0,\infty)$.
+We will assume $X_{1},\dots,X_{n}$ in (3) are super-Fréchet.
+Since $X_{1},\dots,X_{n}$ may not have the same distribution, how to choose the distribution of $X$ is not clear.
+A perhaps natural candidate is the generalized mean of the distributions of $X_{1},\dots,X_{n}$.
+For $r\in\mathbb{R}\setminus\{0\}$, $n\in\mathbb{N}$, and $\mathbf{w}=(w_{1},\dots,w_{n})\in\Delta_{n}$, the generalized $r$-mean function is defined as
+$M^{\mathbf{w}}_{r}(u_{1},\dots,u_{n})=\left(w_{1}u_{1}^{r}+\dots+w_{n}u_{n}^{r}\right)^{1/r},~{}~{}~{}~{}~{}(u_{1},\dots,u_{n})\in(0,\infty)^{n}.$
+The generalized 00-mean function is the weighted geometric mean, that is, $M^{\mathbf{w}}_{0}(u_{1},\dots,u_{n})=\prod_{i=1}^{n}u_{i}^{w_{i}},$ which is also the limit of $M^{\mathbf{w}}_{r}$ as $r\to 0$.
+A generalized mean of distribution functions is a distribution function.
+In particular, if $r=1$, it leads to a distribution mixture model, i.e., if $X\sim M^{\mathbf{w}}_{1}(F_{1},\dots,F_{n})$, $X$ has the same distribution as $\sum_{i=1}^{n}X_{i}\mathds{1}_{A_{i}}$ where $A_{1},\dots,A_{n}$ are mutually exclusive, independent of $X_{1},\dots,X_{n}$, and $\mathbb{P}(A_{i})=w_{i}$ for all $i\in[n]$.
+Theorem 2.
+If $X_{1},\dots,X_{n}$ are super-Fréchet and NLOD with $X_{i}\sim F_{i}$, $i\in[n]$, and $X\sim M^{\bar{\theta}}_{r}(F_{1},\dots,F_{n})$ for some $r\geq 0$, then for $\bar{\theta}\in\Delta_{n}^{+}$,
+$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$
+Proof.
+Let $g_{i}(x)=1/(-\log F_{i}(x))$, $x>0$, for all $i\in[n]$.
+As $g_{i}$, $i\in[n]$, is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g_{i}(x)=0$, $g_{i}(x)\geq\theta g_{i}(x/\theta)$ for all $x>0$ and $\theta\in(0,1)$.
+Then, for $\theta\in(0,1)$,
+$F_{i}\left(\frac{x}{\theta}\right)=\exp\left(-g_{i}\left(\frac{x}{\theta}\right)^{-1}\right)\leq\exp\left(-\theta g_{i}(x)^{-1}\right)=F_{i}(x)^{\theta}.$ (4)
+As $X_{1},\dots,X_{n}$ are NLOD, by (4), for any $x>0$, $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$, and $r\geq 0$,
+$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F_{i}\left(\frac{x}{\theta_{i}}\right)\leq\prod_{i=1}^{n}F_{i}\left(x\right)^{\theta_{i}}$
+$=M^{\bar{\theta}}_{0}(F_{1}(x),\dots,F_{n}(x))\leq M^{\bar{\theta}}_{r}(F_{1}(x),\dots,F_{n}(x))=\mathbb{P}(X\leq x).$
+The last inequality is because the generalized mean function is monotone in $r$; that is, given any $\mathbf{w}\in\Delta_{n}$, $M^{\mathbf{w}}_{r}\leq M^{\mathbf{w}}_{s}$ for $r\leq s$ (Theorem 16 of Hardy et al.
+(1934)).
+∎
+5 Comparison with existing results
+In this section, we compare our results with the literature.
+We first consider the case when $X_{1},\dots,X_{n}$ in (SD) are iid.
+In Arab et al., (2024), it is shown that (SD) holds for non-negative random variables that are InvSub; a random variable $X\sim F$ and its distribution is called InvSub if $1-F(1/x)$ is subadditive.
+The class of InvSub distributions is larger than $\mathcal{H}$ as $h_{F}=-\log F(1/x)$ is subadditive implies that $1-F(1/x)$ is subadditive.
+Müller, (2024) showed that (SD) holds for super-Cauchy random variables; a random variable $X\sim F$ and its distribution is called super-Cauchy if $F^{-1}(G(x))$ is convex where $G$ is the standard Cauchy distribution function.
+Super-Cauchy distributions are continuous and can take positive values on the entire real line but they do not contain $\mathcal{H}$ as $\mathcal{H}$ includes non-continuous distributions (see Example 3).
+The proofs in both Arab et al., (2024) and Müller, (2024) are short and elegant.
+As our results cover the case of negatively dependent risks, for the rest of this section, we will focus on the comparison of our results with Chen et al.
+(2024a); to our best knowledge, Chen et al.
+(2024a) is the only other paper that deals with negatively dependent risks.
+Chen et al.
+(2024a) showed that
+(SD) holds for super-Pareto risks X1,,Xn that are weakly negatively associated.italic-(SDitalic-) $\mbox{~{}holds for super-Pareto risks $X_{1},\dots,X_{n}$ that are weakly negatively associated.}$.
+(5)
+For ease of comparison, definitions of super-Pareto distributions and weak negative association are given in a slightly different form from Chen et al. (2024a) below.
+Definition 5.
+A random variable $X$ and its distribution is super-Pareto if $X$ and $f(Y)$ have the same distribution for some non-decreasing, convex, and non-constant function $f$ with $f(0)=0$ and $Y\sim\mathrm{Pareto}(1)$.
+Definition 6.
+A set $S\subseteq\mathbb{R}^{k}$, $k\in\mathbb{N}$ is decreasing if $\mathbf{x}\in S$ implies $\mathbf{y}\in S$ for all $\mathbf{y}\leq\mathbf{x}$.
+Random variables $X_{1},\dots,X_{n}$ are weakly negatively associated if for any $i\in[n]$, decreasing set $S\subseteq\mathbb{R}^{n-1}$, and $x\in\mathbb{R}$ with $\mathbb{P}(X_{i}\leq x)>0$,
+$\mathbb{P}(\mathbf{X}_{-i}\in S,X_{i}\leq x)\leq\mathbb{P}(\mathbf{X}_{-i}\in S)\mathbb{P}(X_{i}\leq x).$
+where $\mathbf{X}_{-i}=(X_{1},\dots,X_{i-1},X_{i+1},\dots,X_{n})$.
+Lemma 1.
+If random variables $X_{1},\dots,X_{n}$ are super-Pareto and weakly negatively associated, then $X_{1},\dots,X_{n}\in\mathcal{H}$ and they are NLOD.
+Proof.
+As Pareto$(1)$ risks are in $\mathcal{H}$ (see Example 2), by Proposition 2 (iv), super-Pareto risks are in $\mathcal{H}$.
+Since $X_{1},\dots,X_{n}$ are weakly negatively associated, for any $(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$,
+$\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\leq\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n-1}\leq x_{n-1})\mathbb{P}(X_{n}\leq n)\leq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i}).$
+Thus, $X_{1},\dots,X_{n}$ are NLOD. ∎
+The above lemma shows that Theorem 1 implies (5), which is in Theorem 1 (i) of Chen et al.
+(2024a).
+We present below a corollary, which leads to a similar result as Theorem 1 (ii) of Chen et al.
+(2024a).
+Corollary 1.
+Suppose that a random variable $X\in\mathcal{H}$, random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, and $\xi_{1},\dots,\xi_{n}$ are any positive random variables independent of $X,X_{1},\dots,X_{n}$ with $\sum_{i=1}^{n}\xi_{i}\leq 1$.
+If $\mathbb{P}(cX>t)\geq c\mathbb{P}(X>t)$ for all $c\in(0,1]$ and $t>0$, then for $x\geq 0$,
+$\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}(X>x).$ (6)
+Proof.
+By Theorem 1 and the independence between $\xi_{1},\dots,\xi_{n}$ and $X,X_{1},\dots,X_{n}$, we have
+$\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)$ $=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right]$
+$\geq\mathbb{E}\left[\mathbb{P}\left(\left(\sum_{i=1}^{n}\xi_{i}\right)X>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}\left(X>x\right).\qed$
+For $\bar{\theta}\in\Delta_{n}$, let $A_{1},\dots,A_{n}$ be any events independent of $(X_{1},\dots,X_{n})$ and event $A$ be independent of $X$ satisfying $\mathbb{P}(A)=\sum_{i=1}^{n}\theta_{i}\mathbb{P}(A_{i})$.
+If $X_{1},\dots,X_{n}$ are financial losses, $A_{1},\dots,A_{n}$ can be interpreted as the triggering events for these losses.
+Let $\xi_{i}=\theta_{i}\mathds{1}_{A_{i}}$.
+By (6), for $x\geq 0$,
+$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\theta_{i}\mathds{1}_{A_{i}}\right)\mathbb{P}(X>x)=\mathbb{P}(A)\mathbb{P}(X>x)=\mathbb{P}(X\mathds{1}_{A}>x),$
+which is equivalent to
+$X\mathds{1}_{A}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}.$ (7)
+Theorem 1 (ii) of Chen et al. (2024a) showed (7) with different assumptions from Corollary 1; we refer readers to Chen et al. (2024a) for more details.
+6 Conclusion
+In this paper, we provide some sufficient conditions for property (SD) to hold.
+One can see that the property, while very strong, holds for a remarkably large class of distributions.
+We have also shown that it remains valid for non-identically distributed random variables.
+We conclude with some open questions.
+First, we are interested in understanding how close our sufficient conditions for (SD) are to the optimal ones, i.e., we would like to understand what conditions are necessary for (SD).
+Second, the definition of our class of heavy-tailed random variables seems to suggest that it is the distribution of $1/X$ that is of importance. We currently lack an intuitive explanation of this.
+Finally, property (SD) raises the possibility that, for some random variables $X_{1},\dots,X_{n}$ and two vectors $\bar{\eta},\bar{\gamma}\in\mathbb{R}^{n}_{+}$,
+$\eta_{1}X_{1}+\dots+\eta_{n}X_{n}\leq_{\rm st}\gamma_{1}X_{1}+\dots+\gamma_{n}X_{n},$ (8)
+where $\bar{\gamma}$ is smaller than $\bar{\eta}$ in majorization order; that is, $\sum^{n}_{i=1}\gamma_{i}=\sum^{n}_{i=1}\eta_{i}$ and $\sum^{k}_{i=1}\gamma_{(i)}\geq\sum^{k}_{i=1}\eta_{(i)}\,$ for $k\in[n-1]$ where $\gamma_{(i)}$ and $\eta_{(i)}$ represent the $i$th smallest order statistics of $\bar{\gamma}$ and $\bar{\eta}$.
+Clearly, (8) implies (SD).
+It is well known that (8) holds for iid stable random variables with infinite mean (see Ibragimov, (2005)) and it was recently shown to hold for iid Pareto random variables with infinite mean by Chen et al., (2025).
+It is of question whether (8) can hold for a larger class of distributions.
+Note that the methods used in the current paper do not appear to be useful to address (8) as we rely on the comparison of a sum with each of the summands.
+A more subtle approach to sums is required.
+Appendices
+Appendix A Proof of Proposition 2
   1. (i)
-    $F$ is continuous on $[0,\infty)$.
-  2. (ii)
-    $F^{\beta}$, $\beta>0$, is super heavy-tailed.
-  3. (iii)
-    If, in addition, a random variable $Y\sim G$ is super heavy-tailed and independent of $X$, then $\max\{X,Y\}$ is also super heavy-tailed.
-In terms of distribution functions, if $F$ and $G$ are both super heavy-tailed, then so is $FG$.
-  4. (iv)
-    For a non-decreasing, convex, and non-constant function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ with $f(0)=0$, $f(X)$ is super heavy-tailed.
-Proof.
-  1. (i)
-    As $h_{F}$ is sub-additive and increasing, and $\lim_{x\downarrow 0}h_{F}(x)=0$, $h_{F}$ is continuous on $(0,\infty)$, and so is $F$ (see Remark 1 of Matkowski and Świątkowski, (1993)).
+    As $h_{F}$ is subadditive and increasing, and $\lim_{x\downarrow 0}h_{F}(x)=0$, $h_{F}$ is continuous on $(0,\infty)$, and so is $F$ (see Remark 1 of Matkowski and Świątkowski, (1993)).
 The desired result is due to the right-continuity of $F$.
@@
     Hence, we have
-    $\left(f^{-1+}\left(\frac{ab}{a+b}\right)\right)^{-1}\leq\left(f^{-1+}\left(a\right)\right)^{-1}+\left(f^{-1+}\left(b\right)\right)^{-1}.$ (3)
+    $\left(f^{-1+}\left(\frac{ab}{a+b}\right)\right)^{-1}\leq\left(f^{-1+}\left(a\right)\right)^{-1}+\left(f^{-1+}\left(b\right)\right)^{-1}.$ (A.1)
     Denote by $F$ and $G$ the distribution functions of $X$ and $f(X)$, respectively.
@@
 By letting $g(x)=1/f^{-1+}(1/x)$ for $x>0$, we write $h_{G}=h_{F}\circ g$.
-By inequality (3), for any $x,y>0$,
+By inequality (A.1), for any $x,y>0$,
     $g(x+y)=\left(f^{-1+}\left(\frac{1/xy}{1/x+1/y}\right)\right)^{-1}\leq\left(f^{-1+}\left(\frac{1}{x}\right)\right)^{-1}+\left(f^{-1+}\left(\frac{1}{y}\right)\right)^{-1}=g(x)+g(y).$
-    Therefore, $g$ is sub-additive. As $h_{F}$ is sub-additive and non-decreasing, it is clear that $h_{G}=h_{F}\circ g$ is sub-additive and we have the desired result. ∎
-Proposition 3.
-Let $\bar{\theta}\in\Delta_{n}$.
-If distribution functions $F_{1},\dots,F_{n}$ are super heavy-tailed and $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{n}$, then $\sum_{i=1}^{n}\theta_{i}F_{i}$ is also super heavy-tailed.
-Proof.
+    Therefore, $g$ is subadditive. As $h_{F}$ is subadditive and non-decreasing, it is clear that $h_{G}=h_{F}\circ g$ is subadditive and we have the desired result.
+Appendix B Proof of Proposition 3
 Let $G=\sum_{i=1}^{n}\theta_{i}F_{i}$. It suffices to show
-$G\left(\frac{xy}{x+y}\right)\geq G(x)G(y)~{}~{}~{}\mbox{for all $x,y>0$.}$ (4)
+$G\left(\frac{xy}{x+y}\right)\geq G(x)G(y)~{}~{}~{}\mbox{for all $x,y>0$.}$ (A.2)
 For $n=2$, as $F_{1}$ and $F_{2}$ are super heavy-tailed,
@@
 $=\theta_{1}\theta_{2}(F_{1}\left(x\right)-F_{2}(x))(F_{1}\left(y\right)-F_{2}(y))\geq 0.$
-Hence, (4) holds for $n=2$.
-Next, assume that (4) holds for $n=k-1$ where $k>3$ is an integer.
+Hence, (A.2) holds for $n=2$.
+Next, assume that (A.2) holds for $n=k-1$ where $k>3$ is an integer.
 Let $a=\sum_{i=1}^{k-1}\theta_{i}F_{i}(x)$, $b=\sum_{i=1}^{k-1}\theta_{i}F_{i}(y)$, $c=a/(F_{n}(x)(1-\theta_{n}))$, and $d=b/(F_{n}(y)(1-\theta_{n}))$.
@@
 $=\theta_{n}(1-\theta_{n})F_{n}(x)F_{n}(y)\left(cd+1-c-d\right).$
-As $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{k}$, $F_{k}\leq\sum_{i=1}^{k-1}\theta_{i}/(1-\theta_{n})F_{i}$. Thus $c,d\geq 1$ and $cd+1-c-d\geq 0$. The proof is complete by induction. ∎
-Remark 2.
-Let $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ be some function.
-It is well-known that if $f(x)/x$ is decreasing, $f$ is sub-additive (Theorem 7.2.4 of Hille and Phillips, (1996)).
-Another sufficient condition for $f$ being sub-additive is that $f$ is concave and $\lim_{x\downarrow 0}f(x)\geq 0$ (Theorem 7.2.5 of Hille and Phillips, (1996)).
-Hence, if $h_{F}(x)/x$ is decreasing or $h_{F}$ is concave, $F$ is super heavy-tailed.
-3.2 Examples of super heavy-tailed distributions
-In this section, we demonstrate that many well-known infinite-mean distributions belong to the super heavy-tailed class.
-Example 1 (Pareto distribution).
+As $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{k}$, $F_{k}\leq\sum_{i=1}^{k-1}\theta_{i}/(1-\theta_{n})F_{i}$. Thus $c,d\geq 1$ and $cd+1-c-d\geq 0$. The proof is complete by induction.
+Appendix C Examples of distributions in the class $\mathcal{H}$
+In this section, we demonstrate that many well-known infinite-mean distributions are in $\mathcal{H}$.
+Example A.1 (Pareto distribution).
 For $\alpha>0$, the Pareto distribution, denoted by Pareto$(\alpha)$, is defined as
-$\mathbb{P}(X\leq x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
-If $\alpha\leq 1$, the Pareto distribution is super heavy-tailed.
-This can be seen by noting that any Pareto$(\alpha)$ random variable $X$ can be written as $X=f(Z)$ where $Z\sim{\rm Pareto}(1)$ and $f(x)=(x+1)^{1/\alpha}-1$ for $x\geq 0$.
-It is easy to show that Pareto$(1)$ is super heavy-tailed.
-As $f$ is increasing and convex if $\alpha\leq 1$, the result follows from Proposition 2 (iv).
-Example 2 (Fréchet distribution).
-For $\alpha>0$, the Fréchet distribution is defined as
-$\mathbb{P}(X\leq x)=\exp(-x^{-\alpha}),~{}~{}x>0.$
-If $\alpha\leq 1$, the Fréchet distribution is super heavy-tailed.
-To see this, let $F$ denote the distribution function of $X$.
-We have $h_{F}(x)=x^{\alpha}$ for $x>0$.
-Then, if $\alpha\leq 1$, for $x,y>0$,
-$\frac{h_{F}(x)+h_{F}(y)}{h_{F}(x+y)}=\left(\frac{x}{x+y}\right)^{\alpha}+\left(1-\frac{x}{x+y}\right)^{\alpha}\geq 1.$
-Thus $h_{F}$ is sub-additive and the Fréchet distribution with $\alpha\leq 1$ is super heavy-tailed.
-Example 3 (Generalized Pareto distribution).
+$F(x)=1-\frac{1}{(x+1)^{\alpha}},~{}~{}x>0.$
+If $\alpha=1$, for $x,y>0$,
+$h_{F}(x+y)-h_{F}(x)-h_{F}(y)=\log(x+y+1)-\log(x+1)-\log(y+1)\leq 0$
+Thus, Pareto$(1)\in\mathcal{H}_{s}$.
+Then we note that any Pareto$(\alpha)$ random variable $X$ can be written as $X=f(Z)$ where $Z\sim{\rm Pareto}(1)$ and $f(x)=(x+1)^{1/\alpha}-1$ for $x\geq 0$.
+By Proposition 2 (iv), as $f$ is increasing and convex for $\alpha\leq 1$, Pareto$(\alpha)\in\mathcal{H}_{s}$ if $\alpha\leq 1$.
+Example A.2 (Generalized Pareto distribution).
 The generalized Pareto distribution with parameters $\xi\in\mathbb{R}$ and $\beta>0$ is defined as
-$\mathbb{P}(X\leq x)=\begin{cases}1-\left(1+\xi\frac{x}{\beta}\right)^{-1/\xi}&x\in D(\xi,\beta)\\ e^{-x/\beta}&x\geq 0,\end{cases}$
-where $D(\xi,\beta)=[0,\infty)$ if $\xi>0$ and $D(\xi,\beta)=[0,-\beta/\xi)$ if $\xi<0$.
+$F(x)=\begin{cases}1-\left(1+\xi\frac{x}{\beta}\right)^{-1/\xi},&\mbox{~{}~{}if$\xi\neq 0$},\\ e^{-x/\beta},&\mbox{~{}~{}if $\xi=0$},\end{cases}$
+where $x\in[0,\infty)$ if $\xi\geq 0$ and $x\in[0,-\beta/\xi)$ if $\xi<0$.
 By the Pickands-Balkema-de Haan Theorem (Balkema and de Haan,, 1974; Pickands,, 1975), the generalized Pareto distributions are the only possible non-degenerate limiting distributions of the excess of random variables beyond a high threshold.
-If $\xi\geq 1$, then the generalized Pareto distribution is super heavy-tailed.
+If $\xi\geq 1$, $F\in\mathcal{H}$.
 This is by Proposition 2 (iv); that is, the generalized Pareto random variables with $\xi\geq 1$ can be obtained from location-scale transforms of Pareto$(1/\xi)$ random variables.
-Example 4 (Super-Pareto distribution).
-Introduced by Chen et al.
-(2024a), a super-Pareto distribution is the distribution of a random variable $f(X+1)$ where $X\sim\rm{Pareto}(1)$ and $f$ is some non-decreasing, convex, and non-constant function.
-By Proposition 2 (iv), a super-Pareto distribution (after some location-scale transform) is super heavy-tailed.
-Example 5 (Burr distribution).
+Example A.3 (Burr distribution).
 For $\alpha,\tau>0$, the Burr distribution is defined as
-$\mathbb{P}(X\leq x)=1-\left(\frac{1}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0.$ (5)
+$F(x)=1-\left(\frac{1}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0.$ (A.3)
 Let $Y\sim\rm{Pareto}(\alpha)$.
 Then $Y^{1/\tau}$ follows a Burr distribution.
-If $\alpha,\tau\leq 1$, the Burr distribution is super-Pareto and hence super heavy-tailed.
+If $\alpha,\tau\leq 1$, the Burr distribution is super-Pareto and hence $F\in\mathcal{H}$.
 Special cases of Burr distributions are the paralogistic ($\alpha=\tau$) and the log-logistic ($\alpha=1$) distributions; see Kleiber and Kotz, (2003) and Klugman et al.
 (2012).
-Example 6 (Inverse Burr distribution).
-Suppose that $Y$ follows the Burr distribution (5). Then $X=1/Y$ follows the inverse Burr distribution
-$\mathbb{P}(X\leq x)=\left(\frac{x^{\tau}}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0,$
+Example A.4 (Inverse Burr distribution).
+Suppose that $Y$ follows the Burr distribution (A.3). Then $X=1/Y$ follows the inverse Burr distribution
+$F(x)=\left(\frac{x^{\tau}}{x^{\tau}+1}\right)^{\alpha},~{}~{}x>0,$
 where $\alpha,\tau>0$.
-Let $F$ denote the inverse Burr distribution function.
-If $\tau\leq 1$, it is easy to check that the second derivative of $h_{F}$ is always negative, and thus $h_{F}$ is sub-additive (see Remark 2).
-Hence $X$ is super heavy-tailed if $\tau\leq 1$.
-Note that the super heavy-tailed property may not always be preserved under the inverse transformation of super heavy-tailed random variables.
-For instance, if $Z$ follows a Fréchet distribution without finite mean, then $1/Z$ follows a Weibull distribution whose mean is always finite and it cannot be super heavy-tailed.
-Example 7 (Log-Pareto distribution).
+If $\tau\leq 1$, it is easy to check that the second derivative of $h_{F}$ is always negative, and thus $h_{F}$ is subadditive.
+Hence $F\in\mathcal{H}$ if $\tau\leq 1$.
+Note that the property of $\mathcal{H}$ may not always be preserved under the inverse transformation.
+For instance, if $Z$ follows a Fréchet distribution without finite mean, then $1/Z$ follows a Weibull distribution whose mean is always finite.
+Example A.5 (Log-Pareto distribution).
 If $Y\sim\mathrm{Pareto}(\alpha)$, $\alpha>0$, then $X=\exp(Y)-1$ has a log-Pareto distribution (see p. 39 in Arnold, (2015)), with distribution function
-$\mathbb{P}(X\leq x)=1-\frac{1}{(\log(x+1)+1)^{\alpha}},~{}~{}x>0.$
-If $\alpha\in(0,1]$, by Proposition 2 (iv), $X$ is super heavy-tailed.
-Example 8 (Stoppa distribution).
+$F(x)=1-\frac{1}{(\log(x+1)+1)^{\alpha}},~{}~{}x>0.$
+If $\alpha\in(0,1]$, by Proposition 2 (iv), $F\in\mathcal{H}$.
+Example A.6 (Stoppa distribution).
 For $\alpha>0$ and $\beta>0$, a (location-shifted) Stoppa distribution can be defined as
-$\mathbb{P}(X\leq x)=\left(1-\frac{1}{(x+1)^{\alpha}}\right)^{\beta},~{}~{}x>0.$
-Since a Stoppa distribution is a power transform of a Pareto distribution, by Proposition 2 (ii), if $\alpha\leq 1$, Stoppa distribution is super heavy-tailed.
+$F(x)=\left(1-\frac{1}{(x+1)^{\alpha}}\right)^{\beta},~{}~{}x>0.$
+Since a Stoppa distribution is a power transform of a Pareto distribution, by Proposition 2 (ii), if $\alpha\leq 1$, $F\in\mathcal{H}$.
 Power transforms have also been used to generalize Burr distributions (see p.
 211 of Kleiber and Kotz, (2003)).
-Example 9 (Mixture of distributions).
-Suppose that $F_{1},\dots,F_{n}$ are Pareto distributions with possibly different tail parameters $0<\alpha_{1},\dots,\alpha_{n}\leq 1$.
-As $F_{1},\dots,F_{n}$ are comparable in stochastic order, by Proposition 3, mixtures of $F_{1},\dots,F_{n}$ are super heavy-tailed.
-Similarly, mixtures of super heavy-tailed Stoppa distributions are also super heavy-tailed.
-Example 10 (Excess of loss).
-For a super heavy-tailed random variable $X$, by Proposition 2 (iv), $(X-m)_{+}$ is also super heavy-tailed where $m>0$.
-Here, $(X-m)_{+}$ can be interpreted as the loss assumed by a (re)insurer who issues excess-of-loss (re)insurance contracts with retention level $m$.
-4 Weighted sums of non-iid random variables
-So far we have focused on property (2) where $X_{1},\dots,X_{n}$ are independent copies of $X$.
-In this section, we explore what happens for $X_{1},\dots,X_{n}$ which are negatively dependent or do not necessarily have the same marginal distribution.
-4.1 Negatively dependent random variables
-The notion of negative dependence below will be used to establish the main result of this section.
-Definition 3 (Block et al., (1982)).
-Random variables $X_{1},\dots,X_{n}$ are negatively lower orthant dependent (NLOD) if for all $x_{1},\dots,x_{n}\in\mathbb{R}$, $\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\leq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i})$.
-Negative lower orthant dependence is implied by several popular notions of negative dependence in the literature, such as negative association (Alam and Saxena (1981) and Joag-Dev and Proschan (1983)), negative orthant dependence (Block et al., (1982)), and negative regression dependence (Lehmann (1966) and Block et al., (1985)).
-It is also implied by a recently introduced notion called weak negative association (Chen et al.
-(2024a)).
-See Chen et al.
-(2024a) for more details about the implication relations of these notions of negative dependence.
-Examples of NLOD random variables include those following normal distributions with non-positive correlations (see Joag-Dev and Proschan (1983)).
-A straightforward observation allows one to prove a version of Theorem 1 for NLOD super heavy-tailed random variables.
-Theorem 2.
-If a random variable $X$ is super heavy-tailed and random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, then for $\bar{\theta}\in\Delta_{n}$,
-$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (6)
-If $X$ is strictly super heavy-tailed, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
-Proof.
-We provide a proof here for completeness but it is a minor variation of the proof of Theorem 1. Let $X\sim F$. We have, for all $x>0$,
-$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
-$=\exp\left(-\sum_{i=1}^{n}h_{F}\left(\frac{\theta_{i}}{x}\right)\right)\leq\exp\left(-h_{F}\left(\sum_{i=1}^{n}\frac{\theta_{i}}{x}\right)\right)=\exp\left(-h_{F}\left(\frac{1}{x}\right)\right)=F(x).$
-The proof is complete. ∎
-Inequality (6) was established in Theorem 1 (i) of Chen et al.
-(2024a) for super-Pareto random variables $X_{1},\dots,X_{n}$ being weakly negatively associated.
-As weakly negatively associated super-Pareto random variables (after location-scale transforms) are super heavy-tailed and NLOD, Theorem 2 generalizes Theorem 1 (i) of Chen et al.
-(2024a).
-We present below a corollary that generalizes Theorem 1 (ii) of Chen et al.
-(2024a).
-Corollary 1.
-Suppose that a random variable $X$ is super heavy-tailed, random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, and $\xi_{1},\dots,\xi_{n}$ are any positive random variables independent of $X,X_{1},\dots,X_{n}$.
-  1. (i)
-    We have $\sum_{i=1}^{n}\xi_{i}X\leq_{\rm st}\sum_{i=1}^{n}\xi_{i}X_{i}$.
-  2. (ii)
-    If, in addition, $\mathbb{P}(cX>t)\geq c\mathbb{P}(X>t)$ for all $c\in(0,1]$ and $t>0$, and $\sum_{i=1}^{n}\xi_{i}\leq 1$, then for all $x\geq 0$,
-    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}(X>x).$ (7)
-Proof.
-  1. (i)
-    By Theorem 2 and the independence between $\xi_{1},\dots,\xi_{n}$ and $X,X_{1},\dots,X_{n}$, for $x\in\mathbb{R}$,
-    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X>x\right).$
-    This implies (i).
-  2. (ii)
-    With the additional assumptions of (ii), we can again write
-    $\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right].$
-    Thanks to (6), the RHS is greater than or equal to
-    $\mathbb{E}\left[\mathbb{P}\left(\left(\sum_{i=1}^{n}\xi_{i}\right)X>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}\left(X>x\right).$
-    The proof is complete. ∎
-Note that for Corollary 1 to hold, no dependence assumptions are needed between $\xi_{1},\dots,\xi_{n}$.
-We explain below how Theorem 1 (ii) of Chen et al.
-(2024a) can be implied by (7).
-For $\bar{\theta}\in\Delta_{n}$, let $A_{1},\dots,A_{n}$ be any events independent of $(X_{1},\dots,X_{n})$ and event $A$ be independent of $X$ satisfying $\mathbb{P}(A)=\sum_{i=1}^{n}\theta_{i}\mathbb{P}(A_{i})$.
-If $X_{1},\dots,X_{n}$ are financial losses, $A_{1},\dots,A_{n}$ can be interpreted as the triggering events for these losses.
-Let $\xi_{i}=\theta_{i}\mathds{1}_{A_{i}}$.
-By (7), for $x\geq 0$,
-$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\theta_{i}\mathds{1}_{A_{i}}\right)\mathbb{P}(X>x)=\mathbb{P}(A)\mathbb{P}(X>x)=\mathbb{P}(X\mathds{1}_{A}>x).$
-Equivalently, we can write the above inequality as $X\mathds{1}_{A}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}$.
-This result was shown in Theorem 1 (ii) of Chen et al.
-(2024a) for weakly negatively associated and regular super-Pareto random variables, which satisfy the assumptions of Corollary 1 (ii); we refer to Chen et al.
-(2024a) for the definition of regular super-Pareto random variables.
-Thus we have the implication.
-The following corollary shows that the stochastic dominance (6) holds strictly for Pareto distributions with infinite mean.
-Corollary 2.
-Suppose that $X,X_{1},\dots,X_{n}\sim{\rm Pareto}(\alpha)$ with $\alpha\leq 1$ and $X_{1},\dots,X_{n}$ are NLOD. Then for $\bar{\theta}\in\Delta_{n}$,
-$X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}.$ (8)
-Proof.
-Let $F={\rm Pareto}(\alpha)$ with $\alpha\leq 1$.
-By Theorem 2, it suffices to show $h_{F}$ is strictly sub-additive.
-Equivalently, we need to show $D(x,y):=F(1/(x+y))-F(1/x)F(1/y)>0$ for $x,y>0$.
-We have $D(x,y)=xy/((x+1)(y+1)(x+y+1))>0$.
-As Pareto random variables with parameter $\alpha<1$ can be obtained from Pareto random variables with $\alpha=1$ by strictly increasing and convex transforms (see Example 1), the desired result follows from Proposition 1 (iv).
-∎
-The strict inequality (8) plays an important role in many optimization problems (see, e.g., risk exchange in Chen et al.
-(2024b) and optimal bundling in Ibragimov and Walden, (2010)).
-We can see from the proof that as many random variables can be obtained by applying strictly increasing and convex transforms to Pareto$(1)$ random variables, (8) holds for a large class of random variables (see Section 3.2).
-A similar proof shows that Fréchet random variables with a tail parameter strictly less than 1 also satisfy (8).
-Remark 3.
-Suppose $X_{ij}$ for all $i\in[m]$ and $j\in[n]$ are super heavy-tailed random variables such that $X_{i1},\dots,X_{in}$ are independent for each $i\in[m]$, and $X_{1j},\dots,X_{mj}$ are NLOD and identically distributed for each $j\in[n]$.
-Let $Y_{i}=\sum_{j=1}^{n}X_{ij}$ for all $i\in[m]$.
-As stochastic order is closed under convolution (e.g., Theorem 1.A.3 of Shaked and Shanthikumar (2007)), for $(\theta_{1},\dots,\theta_{m})\in\Delta_{m}$, $Y_{1}\leq_{\rm st}\sum_{i=1}^{m}\theta_{i}Y_{i}.$
-Remark 4.
-One may expect positive dependence to make larger values of the sum in (6) more likely and thus the sum more likely to stochastically dominate a single random variable.
-We believe that this intuition does not hold due to the very heavy tails of the random variables under consideration.
-It is known, for instance, that very large values of the sum of iid random variables with heavy tails are usually caused by a single random variable taking a large value, while other random variables are moderate.
-If random variables are positively dependent and some of them do not take large values, it makes others more likely to take moderate values too, hence positive dependence hinders large values.
-The situation is of course reversed for negative dependence.
-However, (6) can still hold for Pareto random variables with infinite mean that are positively dependent via some specific Clayton copula (see Chen et al., 2024c ).
-4.2 Non-identical distributions
-In the previous sections, stochastic dominance relations are obtained for random variables with the same marginal distribution.
-We now look at the case when the random variables are not necessarily identically distributed.
-Given non-identically distributed random variables $X_{1},\dots,X_{n}$ and any $\bar{\theta}\in\Delta_{n}$, the question is to study for which random variable $X$ the following property holds
-$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (9)
-To study this problem, we consider the class of super-Fréchet distributions defined below.
-Definition 4.
-A random variable $X$ with $\operatorname*{ess\text{-}inf}X=0$ is said to be super-Fréchet (or has a super-Fréchet distribution) if the function $g:x\mapsto 1/(-\log\mathbb{P}(X\leq x))$ is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g(x)=0$.
-Super-Fréchet random variables can be obtained by applying strictly increasing and convex transforms to Fréchet random variables with tail parameter 1, and therefore are super heavy-tailed.
-As convex transforms make the tail of random variables heavier, super-Fréchet distributions are more heavy-tailed than Fréchet distribution with tail parameter 1, and thus the name.
-Fréchet distributions with infinite mean, as well as many other distributions in the following example, are super-Fréchet.
-Example 11.
-Pareto, Burr, paralogistic, and log-logistic random variables, all with infinite mean, are super-Fréchet distributions.
-Since all these random variables can be obtained by applying strictly increasing and convex transforms to Pareto random variables with tail parameter $1$ (see Example 4), it suffices to show that a Pareto random variable with tail parameter $1$ is super-Fréchet.
-Write the Pareto distribution with tail parameter 1 as
-$F(x)=1-\frac{1}{x+1}=\exp\left(-\frac{1}{g(x)}\right),~{}~{}~{}x>0,$
-where $g(x)=1/\log(1+1/x)$. It is clear that $g$ is strictly increasing and $\lim_{x\downarrow 0}g(x)=0$. We show $g$ is concave on $(0,\infty)$. We have
-$g^{\prime\prime}(x)=\frac{2-(1+2x)\log(1+1/x)}{x^{2}(1+x)^{2}\log^{3}(1+1/x)}.$
-Let $r(x)=\log\left(1/x+1\right)-2/(1+2x)$, $x>0$.
-It is easy to verify that $r$ is strictly decreasing on $(0,\infty)$ and $r(x)$ goes to 00 as $x$ goes to infinity.
-Thus $r(x)>0$ and $g^{\prime\prime}(x)<0$ for $x\in(0,\infty)$.
-We will assume $X_{1},\dots,X_{n}$ in (9) are super-Fréchet.
-Since $X_{1},\dots,X_{n}$ may not have the same distribution, how to choose the distribution of $X$ is not clear.
-A perhaps natural candidate is the generalized mean of the distributions of $X_{1},\dots,X_{n}$.
-For $r\in\mathbb{R}\setminus\{0\}$, $n\in\mathbb{N}$, and $\mathbf{w}=(w_{1},\dots,w_{n})\in\Delta_{n}$, the generalized $r$-mean function is defined as
-$M^{\mathbf{w}}_{r}(u_{1},\dots,u_{n})=\left(w_{1}u_{1}^{r}+\dots+w_{n}u_{n}^{r}\right)^{1/r},~{}~{}~{}~{}~{}(u_{1},\dots,u_{n})\in(0,\infty)^{n}.$
-The generalized 00-mean function is the weighted geometric mean, that is, $M^{\mathbf{w}}_{0}(u_{1},\dots,u_{n})=\prod_{i=1}^{n}u_{i}^{w_{i}},$ which is also the limit of $M^{\mathbf{w}}_{r}$ as $r\to 0$.
-A generalized mean of distribution functions is a distribution function.
-In particular, if $r=1$, it leads to a distribution mixture model.
-Theorem 3.
-Suppose that $X_{1},\dots,X_{n}$ are super-Fréchet, NLOD, and $X_{i}\sim F_{i}$, $i\in[n]$. For $\bar{\theta}\in\Delta_{n}$ and $0\leq t\leq s$,
-$X^{*}_{s}\leq_{\rm st}X^{*}_{t}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i},$
-where $X^{*}_{r}\sim M^{\bar{\theta}}_{r}(F_{1},\dots,F_{n})$, $r\geq 0$.
-Proof.
-Let $g_{i}(x)=1/(-\log F_{i}(x))$, $x>0$, for all $i\in[n]$.
-As $g_{i}$, $i\in[n]$, is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g_{i}(x)=0$, $g_{i}(x)\geq\theta g_{i}(x/\theta)$ for all $x>0$ and $\theta\in(0,1)$.
-Then, for $\theta\in(0,1)$,
-$F_{i}\left(\frac{x}{\theta}\right)=\exp\left(-g_{i}\left(\frac{x}{\theta}\right)^{-1}\right)\leq\exp\left(-\theta g_{i}(x)^{-1}\right)=F_{i}(x)^{\theta}.$ (10)
-As $X_{1},\dots,X_{n}$ are NLOD, by (10), for any $x>0$, $(\theta_{1},\dots,\theta_{n})\in\Delta_{n}$, and $r\geq 0$,
-$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)\leq\prod_{i=1}^{n}F_{i}\left(\frac{x}{\theta_{i}}\right)\leq\prod_{i=1}^{n}F_{i}\left(x\right)^{\theta_{i}}$
-$=M^{\bar{\theta}}_{0}(F_{1}(x),\dots,F_{n}(x))\leq M^{\bar{\theta}}_{r}(F_{1}(x),\dots,F_{n}(x))=\mathbb{P}(X^{*}_{r}\leq x).$
-The last inequality is because the generalized mean function is monotone in $r$; that is, given any $\mathbf{w}\in\Delta_{n}$, $M^{\mathbf{w}}_{r}\leq M^{\mathbf{w}}_{s}$ for $r\leq s$ (Theorem 16 of Hardy et al.
-(1934)).
-∎
-5 Applications
-This section is devoted to some applications of the stochastic dominance properties studied in this paper.
-5.1 Portfolio diversification
-Since the introduction of the modern portfolio theory (Markowitz (1952)), it is well known that diversification can reduce risks of decision makers in finance and insurance sectors, especially when losses have finite means (Samuelson, (1967)).
-However, diversification can also be detrimental in the presence of extremely heavy-tailed losses, as warned by Fama and Miller, (1972), Embrechts et al., (2002), and Ibragimov, (2005).
-We illustrate below the effects of diversification on super heavy-tailed losses.
-Suppose that an agent needs to allocate their exposure $\theta\in\mathbb{R}_{+}$ over NLOD and identically distributed super heavy-tailed losses $X_{1},\dots,X_{n}$ to minimize the risk.
-They use a risk measure to evaluate their risk; a risk measure is a functional $\rho:\mathcal{X}_{\rho}\rightarrow\mathbb{R}$ where $\mathcal{X}_{\rho}$ is a set of random losses.
-The two regulatory risk measures in insurance and finance are Value-at-Risk (VaR) and Expected Shortfall (ES).
-For a random variable $X\sim F$ and $p\in(0,1)$, VaR is defined as the left quantile
-$\mathrm{VaR}_{p}(X)=F^{-1}(p)=\inf\{t\in\mathbb{R}:F(t)\geq p\},$
-and ES is defined as
-$\mathrm{ES}_{p}(X)=\frac{1}{1-p}\int_{p}^{1}\mathrm{VaR}_{u}(X)\mathrm{d}u.$
-Note that for losses with infinite mean, some risk measures such as ES, may give infinite value and thus are not useful for our analysis.
-For the rest of our discussion, we assume $\mathcal{X}_{\rho}$ contains the convex cone generated by $X_{1},\dots,X_{n}$.
-For two random losses $X$ and $Y$, an agent equipped with a risk measure $\rho$ prefers $X$ over $Y$ if $\rho(X)\leq\rho(Y)$.
-All commonly used risk measures, including VaR and ES, are monotone, that is, $\rho(X)\leq\rho(Y)$ if $X\leq_{\rm st}Y$; see Föllmer and Schied (2016) for a detailed discussion on risk measures.
-The result below directly follows from Theorem 2.
-Proposition 4.
-Suppose that $X_{1},\dots,X_{n}$ are identically distributed, NLOD, and super heavy-tailed.
-For $(\theta_{1},\dots,\theta_{n})\in\mathbb{R}_{+}^{n}$ such that $\sum_{i=1}^{n}\theta_{i}=\theta$ and a monotone risk measure $\rho:\mathcal{X}_{\rho}\rightarrow\mathbb{R}$, we have
-$\rho(\theta X_{1})\leq\rho\left(\sum_{i=1}^{n}\theta_{i}X_{i}\right).$ (11)
-Inequality (11) suggests a perhaps surprising implication: The agent should not diversify over super heavy-tailed losses if they want to minimize their risk, which does not depend on their risk preference.
-This is in sharp contrast to the case of iid finite-mean losses, for which diversification is beneficial.
-If $\rho$ is taken to be VaR, then (11) implies the super-additivity of VaR: For all $p\in(0,1)$,
-$\mathrm{VaR}_{p}(\theta_{1}X_{1})+\dots+\mathrm{VaR}_{p}(\theta_{n}X_{n})\leq\mathrm{VaR}_{p}(\theta_{1}X_{1}+\dots+\theta_{n}X_{n}).$ (12)
-Although (12) is not new, the previous results (e.g., Embrechts et al. (2009)) usually hold in some asymptotic sense (i.e., $p$ is close to 1) rather than for all $p\in(0,1)$.
-5.2 An optimal bundling problem
-Bundling is a market strategy to sell goods as collections.
-We complement and generalize the results of the Vickrey auction model studied by Palfrey, (1983) and Ibragimov and Walden, (2010).
-Suppose that there is a single seller who provides $m$ goods to $n$ buyers in the market.
-Let $2^{M}$ be the power set of $[m]$.
-A bundling decision $\beta=\{B_{1},\dots,B_{l}\}$ is a partition of $[m]$ such that
-  1. (a)
-    $B_{s}\neq\emptyset$ for any $s\in[l]$,
-  2. (b)
-    $B_{s}\cap B_{t}=\emptyset$, for any distinct $s,t\in[l]$
-  3. (c)
-    $\bigcup_{s=1}^{l}B_{s}=[m]$.
-For $j\in[n]$, the $j$th buyer’s valuations for goods are represented by $\mathbf{X}_{j}=(X_{1j},\dots,X_{mj})$ where $X_{ij}$ is their valuation for good $i\in[m]$.
-Assume that $X_{1j},\dots,X_{mj}$ are identically distributed and NLOD super heavy-tailed random variables.
-The valuations of buyers are independent of each other, i.e., $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ are independent.
-Each buyer knows their valuations of the goods but not those of the other buyers; the only available information is the distributions of the other buyers’ valuations.
-Assume that each buyer’s valuation for bundles is additive.
-That is, for a bundle $B\in 2^{M}$, its valuation of buyer $j$ is $v_{j}(B)=\sum_{i\in B}X_{ij}$ and for bundling decision $\beta=(B_{1},\dots,B_{l})$, its valuation of buyer $j$ is $v_{j}(\beta)=\sum_{s=1}^{l}v_{j}(B_{s})$.
-Let the seller distribute its goods via Vickrey auctions.
-In this situation, all buyers will simultaneously submit sealed bids for bundles of goods.
-Whoever submits the highest bid will win the goods and pay the seller the second-highest bid.
-For buyer $j\in[n]$, let $\mathbf{x}_{j}=(x_{1j},\dots,x_{mj})\in\mathbb{R}_{+}^{m}$ be their valuations for goods.
-Note that the valuations of the other buyers are unknown to buyer $j$, and thus are random variables.
-That is, to buyer $j$, $v_{j}(B)=\sum_{i\in B}x_{ij}$ and $v_{s}(B)=\sum_{i\in B}X_{is}$ for any $s\in[n]/\{j\}$ and bundle $B\in 2^{M}$.
-Then the expected surplus of buyer $j$ with valuations $\mathbf{x}_{j}$ for bundle $B\in 2^{M}$ is
-$\mathbb{E}(S_{j}(B,\mathbf{x}_{j}))=$ $\mathbb{P}\left(\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)\mathbb{E}\left(v_{j}(B)-\max_{s\in[n]/\{j\}}v_{s}(B)|\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)$
-The expected surplus of buyer $j$ with valuations $\mathbf{x}_{j}$ for bundling decision $\beta=(B_{1},\dots,B_{l})$ is $\mathbb{E}(S_{j}(\beta,\mathbf{x}_{j}))=\sum_{s=1}^{l}\mathbb{E}(S_{j}(B_{s},\mathbf{x}_{j}))$.
-We say buyer $j$ with valuations $\mathbf{x}_{j}$ prefers a bundling decision $\beta_{1}$ to a bundling decision $\beta_{2}$ if $\mathbb{E}(S_{j}(\beta_{1},\mathbf{x}_{j}))\geq\mathbb{E}(S_{j}(\beta_{2},\mathbf{x}_{j}))$.
-Buyers unanimously prefer $\beta_{1}$ to $\beta_{2}$ if, for almost all of their valuations, they prefer $\beta_{1}$ to $\beta_{2}$.
-Proposition 5.
-Suppose that the valuations of goods are identically distributed, NLOD, and super heavy-tailed. Then the buyers unanimously prefer unbundled sales $\{\{1\},\{2\},\dots,\{m\}\}$.
-That is, the buyers prefer separate auctions if their valuations are super heavy-tailed.
-In contrast, the buyers never prefer separate auctions if the valuations are bounded and there are more than two buyers, as shown by Palfrey, (1983).
-If the valuations are iid positive stable random variables with infinite mean, the statement in Proposition 5 also holds (Theorem 4.1 of Ibragimov and Walden, (2010)).
-Thus, our result generalizes that of Ibragimov and Walden, (2010) in both the distributions and dependence of valuations for goods; Banciu and Ødegaard, (2016) studied the pricing of a bundle of products with dependent valuations.
-Proposition 5 can be shown following the proof of Theorem 4.1 of Ibragimov and Walden, (2010).
-For completeness, we provide a proof in the Appendix.
-5.3 A single-period newsvendor problem
-As one of the classical problems in inventory management, the single-period newsvendor problem is to find the optimal level of inventory such that the expected profit of a store is maximized.
-We consider the most basic single-period newsvendor problem for illustration; recent advances in the literature can be found in Qin et al., (2011).
-At the beginning of the period, the customer demand for a single product $D$ is assumed to be a random variable with density $f$ and distribution $F$ and it will be realized at the end of the period.
-A store needs to determine an inventory level $q$ of the products with fixed cost $v$ per unit.
-At the end of the period, if $D\leq q$, then $q-D$ units which are left over are salvaged by the store for revenue of $g<v$ per unit.
-If $D>q$, then $D-q$ units are “lost" sales cost the store $b$ per unit.
-Let the price of the product $p$ be fixed.
-Assume $g=b=0$ for simplicity.
-Then the expected profit of the store is
-$\mathbb{E}(\Pi(D,q))=$ $\int_{0}^{q}(px-vq)f(x)\mathrm{d}x+\int_{q}^{\infty}(pq-vq)f(x)\mathrm{d}x=p\mathbb{E}(\min(D,q))-vq.$ (13)
-It is well known that the optimal inventory level which maximizes (13) is (e.g., Qin et al., (2011))
-$F^{-1}\left(\frac{p-v}{p}\right)=\operatorname*{arg\,max}_{q\in\mathbb{R}}\{\mathbb{E}(\Pi(D,q))\},$
-where $(p-v)/p$ is called the margin ratio.
-Now assume that the store provides a single product to $2$ customers with demands $X_{1}\sim F_{1}$ and $X_{2}\sim F_{2}$.
-The store can either (1) stock the products at 2 locations to serve the customers separately or (2) stock products together in one location (inventory pooling).
-The optimal inventory levels of the two strategies are thus (1) $F_{1}^{-1}((p-v)/p)+F_{2}^{-1}((p-v)/p)$ and (2) $H^{-1}((p-v)/p)$, where $H$ is the distribution of $X_{1}+X_{2}$.
-It has been observed in the literature that inventory pooling can reduce the inventory level in many model assumptions; see Eppen, (1979).
-However, inventory pooling can sometimes increase the inventory level (i.e., $F_{1}^{-1}((p-v)/p)+F_{2}^{-1}((p-v)/p)\leq H^{-1}((p-v)/p)$), and this is called “inventory anomaly" in Yang and Schrage, (2009).
-Proposition 6.
-Suppose that the demands of the two customers are identically distributed, NLOD, and super heavy-tailed. Then inventory pooling always increases the optimal inventory level.
-Proof.
-This is a direct result from Theorem 2 and (1.A.12) of Shaked and Shanthikumar (2007). ∎
-A similar result was shown in Aydın et al., (2012) assuming that the demands follow a regularly varying distribution and the tail probability of their joint distribution is negligible in some sense.
-Their result requires additional constraints on the margin ratio whereas the above proposition holds for any margin ratios between 0 and 1.
-Another close observation is discussed following Theorem 1 of Bimpikis and Markakis, (2016) for iid demands which follow stable distributions with infinite mean.
-6 Conclusion
-In this paper, we provide some sufficient conditions for property (2) to hold.
-One can see that the property, while very strong, holds for a remarkably large class of distributions.
-We have also shown that it remains valid for random variables which are negatively dependent or non-identically distributed.
-We conclude with some open questions which are topics of our ongoing research.
-First, we are interested in understanding how close our sufficient conditions for (2) are to the optimal ones, i.e., we would like to understand what conditions are necessary for (2).
-Second, the definition of super heavy-tailed random variables seems to suggest that it is the distribution of $1/X$ that is of importance. We currently lack an intuitive explanation of this.
-Finally, property (2) raises the possibility that, for some random variables $X_{1},\dots,X_{n}$ and two vectors $\bar{\eta},\bar{\gamma}\in\mathbb{R}^{n}_{+}$,
-$\eta_{1}X_{1}+\dots+\eta_{n}X_{n}\leq_{\rm st}\gamma_{1}X_{1}+\dots+\gamma_{n}X_{n},$ (14)
-where $\bar{\gamma}$ is smaller than $\bar{\eta}$ in majorization order; that is, $\sum^{n}_{i=1}\gamma_{i}=\sum^{n}_{i=1}\eta_{i}$ and $\sum^{k}_{i=1}\gamma_{(i)}\geq\sum^{k}_{i=1}\eta_{(i)}\,$ for $k\in[n-1]$ where $\gamma_{(i)}$ and $\eta_{(i)}$ represent the $i$th smallest order statistics of $\bar{\gamma}$ and $\bar{\eta}$.
-Clearly, (14) implies (2).
-Inequality (14) has been shown in Chen et al., 2024c for iid Pareto random variables with infinite mean.
-A part of our ongoing research is to show that (14) in fact holds for a much larger class of distributions.
-Note that the methods used in the current paper do not appear to be useful to address (14) as we rely on the comparison of a sum with each of the summands.
-A more subtle approach to sums is required.
-Appendices
-Appendix A Proof of Proposition 5
-Suppose $n\geq 2$.
-Let $\mathbf{x}_{j}=(x_{1j},\dots,x_{mj})\in\mathbb{R}_{+}^{m}$ be the valuations of buyer $j\in[n]$ and $v_{j}(B)=\sum_{i\in B}x_{ij}$, $B\in 2^{M}$.
-Let $(X_{1s},\dots,X_{ms})$ be the valuations of buyer $s\in[n]/\{j\}$ and $X_{is}\sim F$, $i\in[m]$.
-For $B\in 2^{M}$, denote by $H$ the distribution function of $v_{s}(B)=\sum_{i\in[B]}X_{is}$.
-Then by independence of $v_{s}(B)$, $s\in[n]/\{j\}$,
-$\mathbb{P}\left(\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)=H\left(v_{j}(B)\right)^{n-1}.$
-Using integration by parts, we get
-$\mathbb{E}\left(\max_{s\in[n]/\{j\}}v_{s}(B)|\max_{s\in[n]/\{j\}}v_{s}(B)<v_{j}(B)\right)$ $=\int_{0}^{v_{j}(B)}x\mathrm{d}\left(\left(\frac{H\left(x\right)}{H\left(v_{j}(B)\right)}\right)^{n-1}\right)$
-$=v_{j}(B)-\int_{0}^{v_{j}(B)}\left(\frac{H\left(x\right)}{H\left(v_{j}(B)\right)}\right)^{n-1}\mathrm{d}x.$
-By Theorem 1, $H(kx)\leq F(x)$, where $k=\mathrm{card}(B)$. Hence the expected surplus for buyer $j$ with valuations $\mathbf{x}_{j}$ is
-$\mathbb{E}(S_{j}(B,\mathbf{x}_{j}))=\int_{0}^{v_{j}(B)}H\left(x\right)^{n-1}\mathrm{d}x=k\int_{0}^{v_{j}(B)/k}H\left(kx\right)^{n-1}\mathrm{d}x,\leq k\int_{0}^{v_{j}(B)/k}F(x)^{n-1}\mathrm{d}x.$
-As $F(x)^{n-1}$ is increasing for $x\geq 0$, $\int_{0}^{y}F(x)^{n-1}\mathrm{d}x$ is a convex function of $y$.
-By Theorem 3.C.1 of Marshall et al.
-(2011), $F(y_{1},\dots,y_{k})=\sum_{t=1}^{k}\int_{0}^{y_{t}}F(x)^{n-1}\mathrm{d}x$, $(y_{1},\dots,y_{k})\in\mathbb{R}_{+}^{k}$, is Schur-convex, which implies $F(\sum_{t=1}^{k}y_{t}/k,\dots,\sum_{t=1}^{k}y_{t}/k)\leq F(y_{1},\dots,y_{k})$.
-Hence,
-$k\int_{0}^{v_{j}(B)/k}F(x)^{n-1}\mathrm{d}x$ $=\sum_{t\in B}\int_{0}^{\sum_{t\in B}x_{tj}/k}F(x)^{n-1}\mathrm{d}x$
-$\leq\sum_{t\in B}\int_{0}^{x_{tj}}F(x)^{n-1}\mathrm{d}x=\sum_{t\in B}\mathbb{E}(S_{j}(\{t\},\mathbf{x}_{j})).$
-Since $\mathbf{x}_{j}$ is arbitrary and the expected surplus for bundling decisions is additive, the proof is done.
+Competing interests
+The authors declare none.
 Acknowledgements
-The authors thank Qihe Tang for the helpful discussions and comments that led to this paper.
+The authors would like to thank two anonymous referees for their helpful comments.
+The authors also thank Qihe Tang and Ruodu Wang for the valuable discussions and comments on a preliminary version of this paper.
+Yuyu Chen is supported by the 2025 Early Career Researcher Grant from the University of Melbourne.